### v1

# There Was Never a Bottleneck in Concept Bottleneck Models

Abstract
Deep learning representations are often difficult to interpret, which can hinder their deployment in sensitive applications. Concept Bottleneck Models (CBMs) have emerged as a promising approach to mitigate this issue by learning representations that support target task performance while ensuring that each component predicts a concrete concept from a predefined set. In this work, we argue that CBMs do not impose a true bottleneck: the fact that a component can predict a concept does not guarantee that it encodes only information about that concept. This shortcoming raises concerns regarding interpretability and the validity of intervention procedures. To overcome this limitation, we propose Minimal Concept Bottleneck Models (MCBMs), which incorporate an Information Bottleneck (IB) objective to constrain each representation component to retain only the information relevant to its corresponding concept. This IB is implemented via a variational regularization term added to the training loss. As a result, MCBMs support concept-level interventions with theoretical guarantees, remain consistent with Bayesian principles, and offer greater flexibility in key design choices.
1 Introduction
Deep learning (DL) models have become the de facto standard for solving a wide range of tasks, primarily due to their ability to uncover complex patterns in data. However, these patterns often go beyond human comprehension, leading to DL models being perceived as black boxes—their predictions may be accurate, but they are typically difficult to interpret. Despite this, interpretability remains a crucial requirement in several domains, including but not limited to healthcare [1, 2, 3], finance [4, 5], and autonomous driving [6, 7].
Most DL models operate by learning data representations—compressed versions of the input that retain essential information for tackling specific tasks of interest [8]. Concept Bottleneck Models (CBMs) aim to define these representations based on a set human-understandable concepts [9]. Given a task $\bm{y}$ and a set of concepts $\bm{c}=\{c_{j}\}_{j=1}^{m}$, CBMs obtain a representation $\bm{z}$ and are trained to jointly optimize the two following objectives:
  1. (i)
    The task $\bm{y}$ must be perfectly solved.
  2. (ii)
    Each concept $c_{j}$ must be predictable by using only a subset $z_{j}\in\bm{z}$.
As a result of the second objective, CBMs are purported to offer: (i) increased interpretability of the representation space, and (ii) the ability to intervene a concrete concept $c_{j}$ by modifying $z_{j}$ and propagating these changes to the predictions.
However, a phenomenon called information leakage has been observed in CBMs [10, 11], whereby $\bm{z}$ encodes information that is not attributable to the defined concepts. This issue poses two major concerns: (i) it undermines the interpretability of the learned representations, and (ii) it compromises the validity of interventions—modifying $z_{j}$ may alter not only the associated concept $c_{j}$, but also other unintended factors encoded in $z_{j}$. We argue that this issue stems from a fundamental limitation in the current formulation of CBMs: the absence of an explicit Information Bottleneck (IB) [12] that actively constrains $z_{j}$ to exclude information unrelated to $c_{j}$. While the second objective in CBMs encourages each $z_{j}$ to retain $c_{j}$ in its entirety, it does not enforce that $z_{j}$ captures only information about $c_{j}$. In the worst-case scenario, $z_{j}$ could encode the entire input $\bm{x}$ and still satisfy this objective.
We propose a solution to the aforementioned problem by introducing an IB in $z_{j}$ in such a way that it does not only preserve all the information of $c_{j}$, but it also contains information only about $c_{j}$. We call these models Minimal Concept Bottleneck Models (MCBMs), since $z_{j}$ is trained to be a minimal sufficient statistic of $c_{j}$ [13, 14]. This contrasts with traditional CBMs, where $z_{j}$ is optimized to be merely a sufficient statistic of $c_{j}$. In other words, while CBMs simply require that $z_{j}$ preserves all information about $c_{j}$, MCBMs require that $z_{j}$ preserves only all the information about $c_{j}$. We implement the IB constraint via an additional loss term derived using variational approximations. This formulation enables faithful concept interventions without violating Bayesian principles and offers greater flexibility in key choices of the model.
2 Related Work
Information Leakage in Concept Bottleneck Models
This phenomenon has been observed in CBMs trained either sequentially or jointly [10, 11], and it arises when the learned representation $\bm{z}$ encodes information that is not present in the concept set $\bm{c}$. Such leakage negatively impacts both the interpretability and the intervenability of the model. As shown in [15], this issue arises when the Markovian assumption fails—i.e., when the target $\bm{y}$ is not fully determined by the concept set $\bm{c}$, or equivalently, $p(\bm{y}|\bm{c})\neq p(\bm{y}|\bm{c},\bm{x})$, which is typically the case in real-world scenarios. In [16], information leakage is tackled from the perspective of out-of-distribution generalization. Concept Embedding Models (CEMs), introduced in [17], aim to alleviate the trade-off between accuracy and interpretability commonly observed in CBMs. However, we argue that this trade-off is inherently limited by the choice of concept set, and that CEMs may, in fact, be less interpretable than standard CBMs, as they yield more entropic representations—thereby amplifying information leakage. Other CEM-based approaches [18, 19, 20] are also prone to this issue. This critique is further elaborated in [21], which introduces an information-theoretic framework along with experimental benchmarks for evaluating interpretability in the representation space. Almost contemporaneously, [22] proposes a metric for interpretability grounded in the concept of usable mutual information [23].
Information Bottleneck in Representation Learning
The Information Bottleneck (IB) [12] is an information-theoretic framework that aims to balance the trade-off between preserving information about a factor variable and compressing the learned representation. A representation that retains all information about the factor is termed sufficient, while one that contains only information relevant to the factor—i.e., discarding all the irrelevant nuisances—is called minimal [24, 25, 26]. This terminology originates from the theory of minimal sufficient statistics [13, 14]. Direct computation of information-theoretic quantities is often intractable due to the complexity of the representation’s underlying distribution. To address this, variational approximations have been proposed to derive tractable evidence bounds that can be efficiently optimized [27, 28]. In this work, we leverage such approximations to optimize the IB objective with respect to the concepts.
Disentanglement
A closely related topic to CBMs that has been extensively studied is disentanglement. Despite its popularity, there remains no universally agreed-upon definition. However, existing definitions can be broadly categorized into three groups: (i) those in [8, 29, 30, 31, 32], which define a disentangled representation as one where changes in a specific part of the representation correspond to changes in a particular factor of variation, while remaining relatively invariant to other factors; (ii) those from works like [33, 34], which view a disentangled representation as one where modifying a single factor of variation leads to changes in only one component of the representation; and (iii) more recent perspectives [35, 36, 37], which consider a representation disentangled only if it satisfies both of the above properties. Conventional CBMs typically yield representations that align with the second definition. In contrast, MCBMs encourage representations to satisfy both criteria simultaneously by introducing an Information Bottleneck for each concept, thereby aligning with the third—and currently most widely accepted—definition of disentanglement. Our experiments demonstrate that MCBMs produce significantly more disentangled representations.
Table 1: Distributions considered in this work. $f_{\theta}$ is typically modeled by a large neural network while $g^{y}_{\phi}$, $g^{c}_{\phi}$ and $g^{z}_{\phi}$ by simpler neural networks. We consider $\bm{z}$ to be always continuous.
$p_{\theta}(z|x)$ $q_{\phi}(\hat{y}|z)$ $q_{\phi}(\hat{c}_{j}|z_{j})$ $q_{\phi}(\hat{z}_{j}|c_{j})$
Binary - $\text{Bernoulli}\left(g^{y}_{\phi}(z)\right)$ $\text{Bernoulli}\left(g^{c}_{\phi}(z_{j})\right)$
Multiclass - $\text{Categoric}\left(g^{y}_{\phi}(z)\right)$ $\text{Categoric}\left(g^{c}_{\phi}(z_{j})\right)$
Continuous $\mathcal{N}\left(f_{\theta}(x),\sigma^{2}_{x}I\right)$ $\mathcal{N}\left(g^{y}_{\phi}(z),\sigma_{\hat{y}}^{2}I\right)$ $\mathcal{N}\left(g^{c}_{\phi}(z_{j}),\sigma_{\hat{c}}^{2}I\right)$ $\mathcal{N}\left(g^{z}_{\phi}(c_{j}),\sigma_{\hat{z}}^{2}I\right)$
3 An Information-Theoretic Perspective on CBMs
3.1 Data Generative Process
For this scenario, we consider inputs $\bm{x}\in\mathcal{X}$, targets $y\in\mathcal{Y}$, concepts $\bm{c}=\{c_{j}\}_{j=1}^{m}\in\mathcal{C}$ and nuisances $\bm{n}\in\mathcal{N}$, such that $p(\bm{x},\bm{y},\bm{c},\bm{n})=p(\bm{x}|\bm{c},\bm{n})p(\bm{y}|\bm{x})$, i.e., the inputs $\bm{x}$ are described by the concepts $\bm{c}$ and the nuisances $\bm{n}$, and the targets $\bm{y}$ are fully described by the input $\bm{x}$. The only difference between $\bm{c}$ and $\bm{n}$ is that the former are observed, while the latter are not or, in other words, labels on $\bm{c}$ are provided to us. We assume access to a training set $\{\bm{x}^{(i)},\bm{y}^{(i)},\bm{c}^{(i)}\}_{i=1}^{N}$, which defines the empirical distribution $p(\bm{x},\bm{y},\bm{c})=\sum_{i=1}^{N}\delta\left(\bm{x}-\bm{x}^{(i)}\right)\delta\left(\bm{y}-\bm{y}^{(i)}\right)\delta\left(\bm{c}-\bm{c}^{(i)}\right)$. The graphical model corresponding to this generative process is shown in Figure 1(a) for the case of two concepts.
3.2 Vanilla Models
In machine learning, the most commonly studied problem is that of predicting $\bm{y}$ from $\bm{x}$, which serves as a foundation for more specialized tasks. We refer to models trained to address this problem as Vanilla Models (VMs). These models typically operate by first extracting an intermediate representation $\bm{z}$ from the input $\bm{x}$ via an encoder $p_{\theta}(\bm{z}\mid\bm{x})$. Subsequently, a prediction $\hat{y}\in\mathcal{Y}$ is produced from $\bm{z}$ using a task head $q_{\phi}(\hat{\bm{y}}\mid\bm{z})$. Since $\bm{z}$ is intended to facilitate accurate prediction of $\bm{y}$, the mutual information between $\bm{z}$ and $\bm{y}$, denoted $I(Z;Y)$, should be maximized. In Appendix B.1, we formally show that:
$\max_{Z}I(Z;Y)=\max_{\theta,\phi}\mathbb{E}_{p(\bm{x},\bm{y})}\left[\mathbb{E}_{p_{\theta}(\bm{z}|\bm{x})}\left[\log q_{\phi}(\hat{\bm{y}}|\bm{z})\right]\right]$ (1)
Figure 1(b) shows the graphical model of a Vanilla Model with a two-dimensional representation $\bm{z}$. Black edges represent the encoder $p_{\theta}(\bm{z}\mid\bm{x})$, while green edges indicate the task head $q_{\phi}(\hat{\bm{y}}\mid\bm{z})$.
The encoder is typically chosen to be deterministic, i.e., $p_{\theta}(\bm{z}\mid\bm{x})=\delta\left(\bm{z}-f_{\theta}(\bm{x})\right)$. However, for tractability reasons (see Section 4.2), we adopt a stochastic formulation where $p_{\theta}(\bm{z}\mid\bm{x})=\mathcal{N}\left(\bm{z};f_{\theta}(\bm{x}),\sigma_{x}^{2}I\right)$, as summarized in Table 1. The choice of task head $q_{\phi}(\hat{\bm{y}}\mid\bm{z})$ depends on the structure of the output space $\mathcal{Y}$, also detailed in Table 1. The objective in Equation (1) corresponds to minimizing the cross-entropy loss when $\bm{y}$ is binary or multiclass, and to minimizing the mean squared error between $\bm{y}$ and $g^{y}_{\phi}(z)$ when $\bm{y}$ is continuous.
3.3 Concept Bottleneck Models
Vanilla Models generally lack interpretability with respect to the known concepts $\bm{c}$, as the encoder $f_{\theta}$ is often opaque and difficult to analyze. Moreover, these models tend to entangle the concepts in such a way that it becomes intractable to determine how individual concepts influence specific components of the latent representation $\bm{z}$, and consequently the predictions $\hat{y}$. Concept Bottleneck Models (CBMs) have been introduced to address this limitation. In a CBM, each concept $c_{j}$ is predicted from a dedicated latent representation $z_{j}$ via a concept head $q(\hat{c}_{j}\mid z_{j})$. Consequently, $z_{j}$ must encode all information about $c_{j}$—that is, $z_{j}$ must be a sufficient representation for $c_{j}$. This requirement can be formalized as maximizing the mutual information $I(Z_{j};C_{j})$, for which the following identity—proved in Appendix B.2—is employed:
$\max_{Z_{j}}I(Z_{j};C_{j})=\max_{\theta,\phi}\mathbb{E}_{p(\bm{x},c_{j})}\left[\mathbb{E}_{p_{\theta}(z_{j}|\bm{x})}\left[\log q_{\phi}(\hat{c}_{j}|z_{j})\right]\right]$ (2)
As illustrated in Figure 1(c), CBMs extend Vanilla Models by incorporating a concept head $q(\hat{c}_{j}\mid z_{j})$, depicted with blue arrows. These models jointly optimize the objectives in Equations (1) and (2).
As detailed in Table 1, the form of the concept head $q_{\phi}(\hat{c}_{j}\mid z_{j})$ depends on the nature of the concept space $\mathcal{C}$. The objective in Equation (2) corresponds to minimizing the cross-entropy loss when $c_{j}$ is binary or multiclass, and the mean squared error between $c_{j}$ and $g_{\phi}^{c}(z_{j})$ when $c_{j}$ is continuous.
(a) Data Generative Process
(b) Vanilla Model (VM)
(c) Concept Bottleneck Model (CBM)
(d) Minimal Concept Bottleneck Model (MCBM)
Figure 1: Graphical models of the different systems described. We consider two concepts and two-dimensional representations. We provide in Appendix A the analogous figure for $m$ concepts and $m$-dimensional representations. Inputs $\bm{x}$ are defined by some concepts $\{c_{j}\}_{j=1}^{m}$ and nuisances $\bm{n}$; and targets $\bm{y}$ are defined by $\bm{x}$ (gray arrows). Vanilla models obtain the representations $\{z_{j}\}_{j=1}^{m}$ from $\bm{x}$ through the encoder $p_{\theta}(\bm{z}|\bm{x})$ (black arrows) and solve the task $\hat{\bm{y}}$ sequentially through the task head $q_{\phi}(\bm{y}|\bm{x})$ (green arrows). Concept Bottleneck Models try to make a prediction $\hat{c}_{j}$ of each concept $c_{j}$ from one representation $z_{j}$ through the concept head $q(\hat{c}_{j}|z_{j})$ (blue arrows). Minimal CBMs (ours) try to make a prediction $z_{j}$ of each representation $z_{j}$ from one concept $c_{j}$ through the representation head $q(z_{j}|c_{j})$ (red arrows).
4 On the Lack of a Bottleneck in CBMs
As explained in Section 1, one of the main advantages commonly attributed to CBMs over vanilla models is their ability to perform concept-level interventions and propagate the effects to the model’s predictions. For instance, consider a scenario in which we aim to estimate $p(\hat{\bm{y}}\mid c_{j}=\alpha,\bm{x})$—that is, the model’s predicted output $\hat{\bm{y}}$ had the concept $c_{j}$ been set to the value $\alpha$. In CBMs, this intervention is typically performed through the latent representation $z_{j}$, according to:
[…]
However, in the original formulation of CBMs (described in Section 3.3), the conditional distribution $p(z_{j}\mid c_{j})$ is not defined—there is no directed path from $c_{j}$ to $z_{j}$ in Figure 1(c). Intuitively, because $z_{j}$ may encode information about $\bm{x}$ beyond $c_{j}$, it cannot be fully determined by $c_{j}$ alone. This leads to a key question: how can interventions be performed in CBMs if $p(z_{j}\mid c_{j})$ is unknown? In what follows, we address this question and argue that current practices often violate basic Bayesian principles, leading to inconsistencies and limiting the architectural flexibility of the model. To resolve this, we introduce Minimal CBMs (MCBMs), a theoretically grounded framework that addresses these issues and enables principled interventions.
4.1 Intervention Mechanisms in Concept Bottleneck Models
As previously discussed, the conditional distribution $p(z_{j}\mid c_{j})$ is not explicitly defined in standard CBMs. To enable a mechanism for performing interventions, two constraints are typically imposed:
  1. (i)
    Concepts $c_{j}$, for $j=1,2,\dots,m$, are assumed to be binary. If a concept is originally multiclass with $k$ categories, it is converted into $k$ binary concepts (One-vs-Rest [38]).
  2. (ii)
    The concept head is defined as $q_{\phi}(c_{j}\mid z_{j})=\sigma(z_{j})$, where $\sigma$ denotes the sigmoid function.
Given that $\sigma$ is an invertible function, this setup permits defining $p(z_{j}\mid c_{j})=\sigma^{-1}(c_{j})$. However, since $\sigma^{-1}(1)=-\sigma^{-1}(0)=\infty$, this inversion becomes ill-defined at the binary extremes. To address this, in practice, $p(z_{j}\mid c_{j}=0)$ and $p(z_{j}\mid c_{j}=1)$ are approximated as the 5th and 95th percentiles of the empirical distribution of $z_{j}$, respectively [9]. Nevertheless, this practice introduces two issues:
  1. 1.
    One-vs-Rest strategies exhibit the following limitations: (i) individual binary classifiers tend to be biased toward the negative class, (ii) multiple classifiers may simultaneously assign a "positive" label to the same input, and (iii) the predicted probabilities across classifiers are typically uncalibrated [39]. These limitations are further illustrated in Section 5.3.
  2. 2.
    The intervention procedure based on the sigmoid inverse does not adhere to Bayes’ rule, i.e., $p(z_{j}\mid c_{j})=\frac{q_{\phi}(c_{j}\mid z_{j})p(z_{j})}{p(c_{j})}$. While $p(c_{j})$ can be approximated using its empirical distribution, the marginal $p(z_{j})$ remains unknown and is challenging to estimate.
4.2 Minimal Concept Bottleneck Models
As previously discussed, CBMs lack an explicit mechanism for enforcing a bottleneck, which often undermines their primary advantages. To address this limitation, we introduce Minimal Concept Bottleneck Models (MCBMs), which explicitly impose an Information Bottleneck. This ensures that each representation $z_{j}$ not only retains all information about the associated concept $c_{j}$, but also contains information only about $c_{j}$. In other words, $z_{j}$ becomes a minimal sufficient representation of $c_{j}$. To achieve this, we introduce a representation head $q_{\phi}(\hat{z}j\mid c_{j})$ that predicts $z_{j}$ from $c_{j}$. This encourages $z_{j}$ to discard information unrelated to $c_{j}$, as doing so improves the predictive accuracy of $q_{\phi}(\hat{z}_{j}\mid c_{j})$. Formally, this objective corresponds to minimizing the conditional mutual information $I(Z_{j};X\mid C_{j})$. We note that if $z_{j}$ is a representation of $\bm{x}$, then the following propositions are equivalent: (i) $I(Z_{j};X|C_{j})=0$, (ii) the Markov Chain $X\leftrightarrow C_{j}\leftrightarrow Z_{j}$ is satisfied and (iii) $p(z_{j}|c_{j})=p(z_{j}|x)$. To minimize $I(Z_{j};X\mid C_{j})$, we leverage the following identity, which is proven in Appendix B.3:
$\min_{Z_{j}}I(Z_{j};X|C_{j})=\min_{\theta,\phi}\mathbb{E}_{p(x,c_{j})}\left[D_{KL}\left(p_{\theta}(z_{j}|x)||q_{\phi}(\hat{z}_{j}|c_{j})\right)\right]$ (4)
Figure 1(d) illustrates how MCBMs extend CBMs by introducing the representation head $q(\hat{z}_{j}\mid c_{j})$, depicted with red arrows. These models are trained to jointly optimize the objectives given in Equations (1), (2), and (4). Although the KL Divergence in Equation (4) does not admit a closed-form solution in general, we show in Appendix B.4 that, under the distributional assumptions listed in Table 1, it reduces to the mean squared error between $f_{\theta}(x)$ and $g^{z}_{\phi}(c_{j})$.
4.3 Practical Considerations for Optimizing MCBMs
To optimize MCBMs, we combine the three previously introduced objectives, incorporating two key considerations. First, we approximate the expectations over $p(\bm{x},\bm{y})$ and $p(\bm{x},c_{j})$ using the empirical data distribution described in Section 3.1, replacing integrals with summations over the dataset. Second, to enable gradient-based optimization through the stochastic encoder, we apply the reparameterization trick [40], allowing for low-variance gradient estimates during backpropagation. $\mathbb{E}_{p_{\theta}(\bm{z}|\bm{x})}\left[\log q_{\phi}(\bm{y}|\bm{z})\right]\approx\sum_{i}\log q_{\phi}\left(\bm{y}|f^{\prime}_{\theta}\left(x,\epsilon^{(i)}\right)\right)$ and $E_{p_{\theta}(z_{j}|\bm{x})}\left[\log q_{\phi}(\hat{c}_{j}|z_{j})\right]\approx\sum_{i}\log q_{\phi}\left(\hat{c}_{j}\big{|}f^{\prime}_{\theta,j}\left(\bm{x},\epsilon^{(i)}\right)\right)$, where $f^{\prime}_{\theta}(\bm{x},\epsilon)=f_{\theta}(\bm{x})+\sigma^{2}_{x}I\epsilon$ (due to the choice of $p_{\theta}(\bm{z}|\bm{x})$ in Table 1), $\epsilon\sim\mathcal{N}\left(0,I\right)$ and $f^{\prime}_{\theta,j}\left(x,\epsilon\right)$ corresponds to the element $j$ of $f^{\prime}_{\theta}\left(\bm{x},\epsilon\right)$. Combining the considerations above yields the final objective for training MCBMs, as shown in Equation (5), where $\beta$ and $\gamma$ are hyperparameters. The first term corresponds to the objective used in Vanilla Models, the second term is introduced in CBMs, and the third is specific to MCBMs. s. Detailed training algorithms for the various cases listed in Table 1 are provided in Appendix C.
$\begin{split}\max_{\theta,\phi}\sum_{k=1}^{N}\sum_{i}\log q_{\phi}\left(\hat{\bm{y}}\big{|}f^{\prime}_{\theta}\left(x^{(k)},\epsilon^{(i)}\right)\right)+&\beta\sum_{j=1}^{n}\log q_{\phi}\left(\hat{c}_{j}\big{|}f^{\prime}_{\theta,j}\left(\bm{x}^{(k)},\epsilon^{(i)}\right)\right)\\ &-\gamma\sum_{j=1}^{n}D_{KL}\left(p_{\theta}\left(z_{j}|\bm{x}^{(k)}\right)\big{|}\big{|}\ q_{\phi}\left(\hat{z}_{j}|c_{j}^{(k)}\right)\right)\end{split}$ (5)
4.4 Intervention Mechanisms in Minimal Concept Bottleneck Models
As discussed at the beginning of this section, intervention procedures require access to $p(z_{j}\mid c_{j})$. As outlined in Section 4.1, this dependence anchors the intervention process to specific architectural design choices in CBMs. Moreover, such procedures violate Bayesian principles, as the prior distribution $p(z_{j})$ is typically unknown. In MCBMs, since $z_{j}$ is explicitly designed to contain information only about $c_{j}$, modifying $z_{j}$ corresponds to intervening solely on $c_{j}$. This is reflected in the graphical model illustrated in Figure 1(d), where MCBMs introduce a directed path from $c_{j}$ to $z_{j}$ through the intermediate variable $\hat{z}_{j}$. This structure permits the computation of:
[…]
which simplifies to $p(z_{j}\mid c_{j})=q_{\phi}(z_{j}\mid c_{j})$ when $p(z_{j}\mid\hat{z}_{j})=\delta(z_{j}-\hat{z}_{j})$, which holds when the objective in Equation (4) is optimized to convergence—or, equivalently, when $z_{j}$ contains information exclusively about $c_{j}$. As shown in Table 1, we define the encoder distribution as $q_{\phi}(z_{j}\mid c_{j})=\mathcal{N}(g^{z}_{\phi}(c_{j}),\sigma_{\hat{z}}^{2}I)$. The mean function $g^{z}_{\phi}(c_{j})$ is chosen according to the following rules:
  1. (i)
    For binary concepts, $g^{z}_{\phi}(c_{j})=\lambda$ if $c_{j}=1$, and $g^{z}_{\phi}(c_{j})=-\lambda$ otherwise.
  2. (ii)
    For categorical concepts, $g^{z}_{\phi}(c_{j})=\lambda\cdot\mathrm{one\_hot}(c_{j})$. This mirrors the Prototypical Learning [41] approach where class-dependent prototypes $\{g^{z}_{\phi}(c_{j})\}_{j=1}^{m}$ are fixed, as adopted in [42].
  3. (iii)
    For continuous concepts, $g^{z}_{\phi}(c_{j})=\lambda\cdot c_{j}$.
Here, $\lambda$ is a scaling constant that controls the norm of the latent representation and is fixed to $\lambda=3$ across all experiments. Regarding the variance term, we set: (i) $\sigma_{x}=0$in the case of CBMs to obtain a deterministic encoder in line with their original formulation, and (ii) $\sigma_{x}=\sigma_{\hat{z}}=1$for MCBMs.
5 Experiments
In this section, we present a series of experiments designed to empirically demonstrate that CBMs fail to impose an effective bottleneck, even in simple settings. We also examine the consequences of this limitation. In contrast, we show that MCBMs successfully enforce a bottleneck, thereby mitigating the issues that arise in its absence. Additional implementation details, including encoder architectures and training hyperparameters for each experiment, are provided in Appendix D.
5.1 Do CBMs and MCBMs Leak Information?
(a) Data
(b) Representation
Figure 2: Some nuisances $\bm{n}_{y}\in\bm{n}$ affect the task $\bm{y}$ while others $\bm{n}_{\bar{y}}\in n$ do not. None of them should affect the representation $\bm{z}$ since it must be fully described by the concepts $\bm{c}$.
In some works, it is assumed that the task $\bm{y}$ is fully defined by the concepts $\bm{c}$. However, it is unrealistic to have a finite set of concepts that we understand as humans and that totally describe an arbitrarily complex task. In general, some of the nuisances $\bm{n}$ that describe the input $\bm{x}$ will also influence $\bm{y}$. Specifically, we decompose the nuisances as $\bm{n}=\{\bm{n}_{y},\bm{n}_{\bar{y}}\}$, where $\bm{n}_{y}$ captures nuisances that, together with the concepts, describe $\bm{y}$, while $\bm{n}_{\bar{y}}$ comprises those that are independent of $\bm{y}$ (see Figure 2(a)). As shown through this text, CBMs have no incentive to remove information that is not related to the concepts. Additionally, since the objective of models in general is to solve $\bm{y}$, they have incentives to preserve not only $\bm{c}$ but also $\bm{n}_{y}$ in their representations. Although this may lead to improved performance in solving the task, it comes at the expense of interpretability and the ability to intervene in the system. For example if $z_{j}\in\bm{z}$ were intended to represent the concept $c_{j}\in\bm{c}$, then one could believe that an intervention is made solely on $c_{j}$ by modifying $z_{j}$. However $z_{j}$ could also contain information on $\bm{n}_{y}$ (or even $\bm{n}_{\bar{y}}$, as we see later), so we would be intervening also the nuisances by modifying $z_{j}$, thereby invalidating any causal conclusion drawn. By introducing the Information Bottleneck in equation (4), we force the model to remove the nuisances from $z_{j}$, thus restoring the validity of any causal analysis about $c_{j}$. We note that this will necessarily negatively impact the performance in solving the task $\bm{y}$ by the nature of the problem. If $\bm{c}$ is incomplete and information from $\bm{n}$ is necessary to solve $\bm{y}$ (i.e., $\bm{n}_{\bar{y}}\neq\emptyset$), the performance will decrease if $\bm{z}$ does not contain information about $\bm{n}$. We analyze next i $\bm{n}_{y}$ and $\bm{n}_{\bar{y}}$ are present in the representation for different versions of CBMs and datasets.
Table 2: Average value of URR for task-related nuisances $\bm{n}_{y}$.
CIFAR-10 MPI3D Shapes3D CelebA CUB
Vanilla $33.5\pm 4.3$ $35.0\pm 1.9$ $45.5\pm 6.4$ $10.7\pm 3.5$ $3.8\pm 1.0$
CBM $34.7\pm 0.2$ $28.1\pm 0.5$ $8.1\pm 2.9$ $11.5\pm 3.5$ $3.8\pm 0.8$
CEM $35.1\pm 0.1$ $43.2\pm 5.2$ $15.8\pm 3.9$ $16.8\pm 3.3$ $3.9\pm 1.1$
MCBM (low $\gamma$) $24.8\pm 0.6$ $10.7\pm 0.1$ $2.5\pm 0.1$ $4.3\pm 0.0$ $3.4\pm 0.9$
MCBM (medium $\gamma$) $19.1\pm 1.0$ $6.7\pm 0.2$ $0.2\pm 0.3$ $\bm{7.9\pm 2.9}$ $2.8\pm 0.8$
MCBM (high $\gamma$) $\bm{11.0\pm 0.5}$ $\bm{0.0\pm 0.0}$ $\bm{0.0\pm 0.0}$ $8.0\pm 2.4$ $\bm{2.4\pm 1.0}$
Task-related information leakage
(a) CBM ($\beta=1$)
(b) MCBM ($\beta=1$, $\gamma=5$)
Figure 3: Average prediction $\hat{y}$ in animal (0) vs. vehicle (1) classification for the different classes when excluding airplane and bird from the concept set.
(c) CBM ($\beta=1$)
(d) MCBM ($\beta=1$, $\gamma=5$)
Figure 4: Average prediction $\hat{y}$ for animal (purple) vs. vehicle (yellow) when intervening on $\bm{z}$ to match the x-axis class instead of the true class (y-axis).
We analyze whether $\bm{n}_{y}$ is encoded in $\bm{z}$—that is, whether the dependency illustrated by the orange arrow in Figure 2(b) is present—and examine its consequences. To illustrate this, we design a simple experiment based on the CIFAR-10 dataset [43]. Since all classes in CIFAR-10 correspond to either animals or vehicles, we define the target $\bm{y}$ as the binary task of distinguishing between these two categories. The concepts $\bm{c}$ are defined using all classes except airplane and bird. Under a proper bottleneck, the model should be unable to accurately predict $\bm{y}$ for these excluded classes, as their information is not captured by any concept. Figure 4 shows the mean predictions $\hat{y}$ for each class, comparing a CBM and an MCBM. The CBM confidently predicts $\bm{y}$ even for the excluded classes, indicating a substantial degree of information leakage. In contrast, the MCBM significantly mitigates this issue, displaying low confidence for excluded classes while maintaining high confidence for those included in the concept set. To assess the impact on interventions, Figure 4 shows model predictions after simulating an intervention corresponding to each concept. In the CBM, predictions following interventions exhibit reduced confidence for certain classes—an effect attributed to residual nuisance information in the representation. This effect is notably reduced in the MCBM.
To systematically analyze this phenomenon, we consider several datasets and define distinct task–concept configurations for each. Specifically: (i) In CIFAR-10 [43], we adopt the described scenario: $\bm{y}$ corresponds to a binary classification between animals and vehicles, $\bm{n}_{y}$ includes airplane and bird, and $\bm{c}$ represents the remaining classes. (ii) In MPI3D [44], $\bm{y}$ corresponds to the object shape, $\bm{n}_{y}$ to the horizontal axis, $\bm{n}_{\bar{y}}$ to the vertical axis, and $\bm{c}$ to the remaining generative factors. (iii) In Shapes3D [30], $\bm{y}$ is the shape, $\bm{n}_{y}$ includes the floor color and wall color, $\bm{n}_{\bar{y}}$ is the orientation, and $\bm{c}$ includes the remaining factors. (iv) In CelebA [45], $\bm{y}$ is gender, $\bm{c}$ consists of eight randomly selected attributes, and $\bm{n}_{y}$ includes the remaining factors, all of which are correlated with $\bm{y}$. (v) In CUB [46], $\bm{y}$ corresponds to the bird species, $\bm{c}$ includes the concepts from seven randomly selected attribute groups, and $\bm{n}_{y}$ comprises the attributes from the remaining 20 groups. Further details for each setup are provided in Appendix D. We recall that, while our MCBMs natively support multiclass concepts, the other models evaluated in this work are limited to binary concepts. Therefore, to ensure fair comparisons, factors with $m$ classes are represented as $m$ separate binary concepts.
To assess the presence of a nuisance factor $n_{j}\in\bm{n}_{y}$ in $\bm{z}$, we aim to estimate $I(N_{j};Z\mid C)$—that is, the amount of additional information that $\bm{z}$ provides about $n_{j}$ beyond what is already captured by $\bm{c}$. Since this quantity is intractable in practice, we approximate it as $\hat{I}(N_{j};Z|C)=\hat{H}(N_{j}|C)-\hat{H}(N_{j}|C,Z)$, where $\hat{H}(N_{j}\mid C)=-\sum_{k=1}^{N}\log h_{\psi}^{c}(\bm{c}^{(k)})$ and $\hat{H}(N_{j}\mid C,Z)=-\sum_{k=1}^{N}\log h_{\psi}^{cz}(\bm{c}^{(k)},\bm{z}^{(k)})$. Here, $h_{\psi}^{c}$ and $h_{\psi}^{cz}$ are MLP classifiers trained to predict $n_{j}$ from $\bm{c}$ and $(\bm{c},\bm{z})$, respectively. In Table 2, we report the average value of $\frac{\hat{I}(N_{j};Z\mid C)}{H(N_{j})}$ across all $n_{j}\in\bm{n}_{y}$, which we refer to as the Uncertainty Reduction Ratio (URR). From these results, we draw the following conclusions: (i) regular CBMs do not consistently remove more nuisance information than VMs—in fact, contrary to common belief, CBMs can sometimes impose a weaker bottleneck than VMs; (ii) CEMs representations are those with more nuisances, i.e., they always introduce a weaker bottleneck even than VMs; and (iii) MCBMs are the models removing more nuisances from the representation, especially when $\gamma$ increases.
Task-unrelated information leakage
Table 3: Average URR for task-unrelated nuisances.
MPI3D Shapes3D
Vanilla $11.3\pm 0.1$ $42.7\pm 9.1$
CBM $7.4\pm 1.9$ $20.6\pm 3.3$
CEM $15.5\pm 4.2$ $40.9\pm 1.8$
MCBM (low $\gamma$) $\bm{0.0\pm 0.0}$ $\bm{0.0\pm 0.0}$
MCBM (medium $\gamma$) $\bm{0.0\pm 0.0}$ $\bm{0.0\pm 0.0}$
MCBM (high $\gamma$) $\bm{0.0\pm 0.0}$ $\bm{0.0\pm 0.0}$
On the one hand, neural networks typically discard information from the input, as their layers are generally non-invertible [47, 48]. On the other hand, there is no incentive to retain $\bm{n}_{\bar{y}}$ in the representation $\bm{z}$, since it is irrelevant for predicting $\bm{y}$. From this, one might expect that $\bm{z}$ should not contain information about $\bm{n}_{\bar{y}}$. However, prior work has shown that neural representations often retain information that is not directly relevant to the target task [24, 49]. To investigate this phenomenon, we analyze if $\bm{n}_{\bar{y}}$ is present in $\bm{z}$ for MPI3D and Shapes3D—the only settings in our experimental setup where $\bm{n}_{\bar{y}}$ is non-empty. We report the URR values for the nuisance variables in $\bm{n}_{\bar{y}}$ in Table 3, and observe the following: (i) in VMs, CBMs, and CEMs, nuisance variables remain present in the representation, even when they are irrelevant to the task $\bm{y}$; and (ii) MCBMs consistently eliminate task-irrelevant nuisances across all values of $\gamma$, which is expected: since there is no incentive to preserve $\bm{n}_{\hat{y}}$, and MCBMs explicitly encourage its removal, such information tends to be discarded.
Are concepts removed to a greater extent in MCBMs?
Table 4: Average concepts accuracy
CelebA CUB
CBM $\bm{92.2\pm 0.7}$ $96.5\pm 0.1$
CEM $92.1\pm 0.8$ $96.5\pm 0.1$
MCBM (low $\gamma$) $92.0\pm 0.8$ $\bm{96.7\pm 0.1}$
MCBM (medium $\gamma$) $92.1\pm 1.0$ $96.4\pm 0.1$
MCBM (high $\gamma$) $91.8\pm 1.1$ $95.8\pm 0.4$
One might suspect that removing nuisance information from the representation could inadvertently eliminate information about the concepts as well. To investigate this, we report concept prediction accuracy for CelebA and CUB in Table 4 (similar trends were observed on the remaining datasets). We observe no significant overall differences across models: while MCBMs show slightly lower concept accuracy on CelebA, they outperform the baselines on CUB for lower $\gamma$. A regular trend, however, is that increasing the hyperparameter $\gamma$ in MCBMs leads to a gradual decrease in concept accuracy, which is consistent with our expectations—stronger regularization may inadvertently suppress features that correlate with the concepts. These results suggest that MCBMs are effective at removing nuisance information while largely preserving the concept-relevant content in the representation.
5.2 Do MCBMs Yield More Interpretable Representations?
Figure 5: CKA (left) and Disentanglement (right).
One of the key properties often attributed to CBMs is that their internal representations align with human-interpretable concepts. While this generally holds, we show that MCBMs yield representations that are even more interpretable. To support this claim, we employ two metrics: (i) Centered Kernel Alignment(CKA) [50, 51, 52] , computed between the learned representations and the concept labels (encoded as one-hot vectors), which quantifies the similarity between representation spaces using kernel methods; and (ii) Disentanglement[36], which assesses whether each dimension of the representation $z_{j}$ captures at most one concept $c_{j}$. We argue that these metrics offer meaningful evidence of interpretability because (i) alignment between the representations and the concepts indicates whether the information is organized in a concept-aware fashion; and (ii) disentanglement reflects the extent to which individual concepts are independently encoded in the learned representations. In Figure 5, we observe that: (i) although the representations learned by CBMs and CEMs exhibit substantially higher alignment with the concepts than VMs, their disentanglement levels are not reliably superior; and (ii) MCBMs consistently achieve improvements in both alignment and disentanglement. Thus, we conclude that removing nuisances from the representations leads to more interpretable representations.
5.3 What Are the Pitfalls of Binarizing Multiclass Concepts?
(a) One-vs-Rest
(b) Multiclass
Figure 6: Class Boundaries
(c) One-vs-Rest
(d) Multiclass
Figure 7: Highest likelihood
(a) One-vs-Rest
(b) Multiclass
Figure 8: Second highest likelihood
As discussed in Section 4.1, CBMs require binarizing multiclass concepts to support interventions—i.e., they adopt a One-vs-Rest classification scheme. In contrast, as discussed in Section 4.4, MCBMs provide a natural and principled framework for modeling and intervening on multiclass concepts directly. Binarizing multiclass variables can introduce several issues, particularly in the presence of class imbalance, which is common in real-world applications. To illustrate this limitation, we design an experiment using a four-class spiral dataset with imbalanced class distributions. We train two CBMs for comparison: (i) one that binarizes concepts using a One-vs-Rest approach, and (ii) one that models concepts directly as multiclass variables.
First, as shown in Figure 8, the One-vs-Rest model tends to favor the most frequent class more often than the Multiclass model, particularly when the true class has low prevalence and is similar (i.e., spatially close in the spiral setting) to the dominant class. Second, Figure 8 illustrates that, in the Multiclass model, the predicted class likelihood increases gradually from regions far from the decision boundaries toward the boundaries, resulting in a well-calibrated probability distribution. Such patterns are absent in the One-vs-Rest model. Finally, in Figure 8, we observe that the second-highest predicted likelihood in the One-vs-Rest model is close to one in certain regions—indicating that multiple classes are simultaneously assigned high confidence. In contrast, the Multiclass model assigns low second-highest likelihoods in regions far from the boundaries, with values increasing only near the decision boundaries.
6 Conclusions
In this paper, we argue that—contrary to common belief—Concept Bottleneck Models (CBMs) do not enforce a true bottleneck: while representations are encouraged to retain concept-related information, they are not constrained to discard nuisance information. This limitation results in a lack of theoretical guarantees for intervention procedures and reduced interpretability in the representation space. To address this, we propose Minimal Concept Bottleneck Models (MCBMs), which introduce an Information Bottleneck in the representation space via an additional loss term derived through variational approximations. Beyond establishing a proper bottleneck, our formulation ensures that interventions do not violate Bayesian principles and introduces greater design flexibility compared to standard CBMs. Through empirical analysis, we show that CBMs—and their existing variants—fail to effectively remove nuisance information from the representation, even when such information is irrelevant to the target task. In contrast, MCBMs successfully eliminate nuisances while preserving concept-relevant information, resulting in more interpretable representations and principled intervention mechanisms. Finally, we highlight a common limitation of CBMs: the need to binarize multiclass concepts due to architectural constraints. MCBMs natively support multiclass concepts, thereby overcoming the drawbacks introduced by this binarization process
Acknowledgments and Disclosure of Funding
This work has received funding from the European Union’s Horizon 2020 research and innovation programme under the Marie Skłodowska-Curie grant agreement No 101007666, MCIN/AEI/10.13039/501100011033 under Grant PID2021-126061OB-C44, and the Government of Aragón (Grant Group T36 23R).
Appendix A Complete Graphical Models
(a) Data Generative Process
(b) Vanilla Model
(c) Concept Bottleneck Models
(d) Minimal Concept Bottleneck Models (ours)
Figure 9: Graphical models of the different systems described. We consider $m$ concepts and $m$-dimensional representations. Inputs $\bm{x}$ are defined by some concepts $\{c_{j}\}_{j=1}^{m}$ and nuisances $\bm{n}$; and targets $\bm{y}$ are defined by $\bm{x}$ (gray arrows). Vanilla models obtain the representations $\{z_{j}\}_{j=1}^{m}$ from $\bm{x}$ through the encoder $p_{\theta}(z|x)$ (black arrows) and solve the task $\hat{y}$ sequentially through the task head $q_{\phi}(y|z)$ (green arrows). Concept Bottleneck Models try to make a prediction $\hat{c}_{j}$ of each concept $c_{j}$ from one representation $z_{j}$ through the concept head $q(\hat{c}_{j}|z_{j})$ (blue arrows). Minimal CBMs (ours) try to make a prediction $\hat{z}_{j}$ of each representation $z_{j}$ from one concept $c_{j}$ through the representation head $q(\hat{z}_{j}|c_{j})$ (red arrows).
Appendix B Details of Derivations
B.1 Proof of Equation 1
$I(Z;Y)$ $=\iint p(z,y)\log\frac{p(y|z)}{p(y)}\,dy\,dz$ (7)
$=\iiint p(x,y)p_{\theta}(z|x)\log\frac{p(y|z)}{p(y)}\,dx\,dy\,dz$ (8)
$=\iiint p(x,y)p_{\theta}(z|x)\log\frac{p(y|z)}{p(y)}\frac{q_{\phi}(\hat{y}|z)}{q_{\phi}(\hat{y}|z)}\,dx\,dy\,dz$ (9)
$=\mathbb{E}_{p(x,y)}\left[\mathbb{E}_{p_{\theta}(z|x)}\left[\log q_{\phi}(\hat{y}|z)\right]\right]+E_{p_{\theta}(z)}\left[D_{KL}\left(p(y|z)||q_{\phi}(\hat{y}|z)\right)\right]+H(Y)$ (10)
$\geq\mathbb{E}_{p(x,y)}\left[\mathbb{E}_{p_{\theta}(z|x)}\left[\log q_{\phi}(\hat{y}|z)\right]\right]+H(Y)$ (11)
Thus, since $H(Y)$ is independent of $\theta$ and $\phi$, we have that:
$\max_{Z}I(Z;Y)=\max_{\theta,\phi}\mathbb{E}_{p(x,y)}\left[\mathbb{E}_{p_{\theta}(z|x)}\left[\log q_{\phi}(\hat{y}|z)\right]\right]$ (12)
B.2 Proof of Equation 2
$I(Z_{j};C_{j})$ $=\iint p(z_{j},c_{j})\log\frac{p(c_{j}|z_{j})}{p(c_{j})}\,dc_{j}\,dz_{j}$ (13)
$=\iiint p(x,c_{j})p_{\theta}(z_{j}|x)\log\frac{p(c_{j}|z_{j})}{p(c_{j})}\,dx\,dc_{j}\,dz_{j}$ (14)
$=\iiint p(x,c_{j})p_{\theta}(z_{j}|x)\log\frac{p(c_{j}|z_{j})}{p(c_{j})}\frac{q(\hat{c}_{j}|z_{j})}{q(\hat{c}_{j}|z_{j})}\,dx\,dc_{j}\,dz_{j}$ (15)
$\geq E_{p(x,c_{j})}\left[E_{p_{\theta}(z_{j}|x)}\left[\log q(\hat{c}_{j}|z_{j})\right]\right]+E_{p_{\theta}(z_{j})}\left[D_{KL}\left(p(c_{j}|z_{j})||q(\hat{c}_{j}|z_{j})\right)\right]+H(C_{j})$ (16)
$\geq E_{p(x,c_{j})}\left[E_{p_{\theta}(z_{j}|x)}\left[\log q(\hat{c}_{j}|z_{j})\right]\right]+H(C_{j})$ (17)
Given the fact that $H(C_{j})$ is constant, we have that:
$\max_{Z_{j}}I(Z_{j};C_{j})=\max_{\theta}E_{p(x,c_{j})}\left[E_{p_{\theta}(z_{j}|x)}\left[\log q(\hat{c}_{j}|z_{j})\right]\right]$ (18)
B.3 Proof of Equation 4
$I(Z_{j};X|C_{j})$ $=\iiint p(x,c_{j})p_{\theta}(z_{j}|x)\log\frac{p(z_{j}|x)}{p(z_{j}|c_{j})}\,dx\,dc_{j}\,dz_{j}$ (19)
$=\iiint p(x,c_{j})p_{\theta}(z_{j}|x)\log\frac{p(z_{j}|x)}{p(z_{j}|c_{j})}\frac{q(\hat{z}_{j}|c_{j})}{q(\hat{z}_{j}|c_{j})}\,dx\,dc_{j}\,dz_{j}$ (20)
$=\mathbb{E}_{p(x,c_{j})}\left[D_{KL}\left(p_{\theta}(z_{j}|x)||q(\hat{z}_{j}|c_{j})\right)\right]-E_{p(c_{j})}\left[D_{KL}\left(p(z_{j}|c_{j})||q(\hat{z}_{j}|c_{j})\right)\right]$ (21)
$\leq\mathbb{E}_{p(x,c_{j})}\left[D_{KL}\left(p_{\theta}(z_{j}|x)||q(\hat{z}_{j}|c_{j})\right)\right]$ (22)
Thus, we have that:
$\min_{Z_{j}}I(Z_{j};X|C_{j})=\min_{\theta}E_{p(x,c_{j})}\left[D_{KL}\left(p_{\theta}(z_{j}|x)||q(\hat{z}_{j}|c_{j})\right)\right]$ (23)
B.4 KL Divergence between two Gaussian Distributions
We are given the conditional distributions:
$p_{\theta}(z_{j}|x)$ $=\mathcal{N}(f_{\theta}(x)_{j},\sigma_{x}^{2}I),$
$q_{\phi}(\hat{z}_{j}|c_{j})$ $=\mathcal{N}(g^{z}_{\phi}(c_{j}),\sigma_{\hat{z}}^{2}I),$
and aim to minimize the expected KL divergence:
$\min_{\theta,\phi}\mathbb{E}_{p(x,c_{j})}\left[D_{\text{KL}}\left(p_{\theta}(z_{j}|x)\;\|\;q_{\phi}(\hat{z}_{j}|c_{j})\right)\right].$
The KL divergence between two multivariate Gaussians with diagonal covariances is given by:
$D_{\text{KL}}(\mathcal{N}(\mu_{p},\Sigma_{p})\;\|\;\mathcal{N}(\mu_{q},\Sigma_{q}))=\frac{1}{2}\left[\operatorname{tr}(\Sigma_{q}^{-1}\Sigma_{p})+(\mu_{q}-\mu_{p})^{\top}\Sigma_{q}^{-1}(\mu_{q}-\mu_{p})-d+\log\frac{\det\Sigma_{q}}{\det\Sigma_{p}}\right].$
Applying this to our case:
  • $\mu_{p}=f_{\theta}(x)_{j}$, $\mu_{q}=g^{z}_{\phi}(c_{j})$,
  • $\Sigma_{p}=\sigma_{x}^{2}I$, $\Sigma_{q}=\sigma_{\hat{z}}^{2}I$,
  • $d$ is the dimension of $z_{j}$.
Plugging in, we obtain:
$D_{\text{KL}}=\frac{1}{2}\left[\frac{d\sigma_{x}^{2}}{\sigma_{\hat{z}}^{2}}+\frac{1}{\sigma_{\hat{z}}^{2}}\|f_{\theta}(x)_{j}-g^{z}_{\phi}(c_{j})\|^{2}-d+d\log\left(\frac{\sigma_{\hat{z}}^{2}}{\sigma_{x}^{2}}\right)\right].$
Note that all terms except the squared distance are constant with respect to $\theta$ and $\phi$. Therefore:
$\mathbb{E}_{p(x,c_{j})}\left[D_{\text{KL}}\left(p_{\theta}(z_{j}|x)\;\|\;q_{\phi}(\hat{z}_{j}|c_{j})\right)\right]=\frac{1}{2\sigma_{\hat{z}}^{2}}\mathbb{E}_{p(x,c_{j})}\left[\|f_{\theta}(x)_{j}-g^{z}_{\phi}(c_{j})\|^{2}\right]+\text{const}.$
Conclusion: Minimizing the expected KL divergence
$\min_{\theta,\phi}\mathbb{E}_{p(x,c_{j})}\left[D_{\text{KL}}\left(p_{\theta}(z_{j}|x)\;\|\;q_{\phi}(\hat{z}_{j}|c_{j})\right)\right]$
is equivalent (up to a scaling factor) to minimizing the expected mean squared error:
$\min_{\theta,\phi}\mathbb{E}_{p(x,c_{j})}\left[\|f_{\theta}(x)_{j}-g^{z}_{\phi}(c_{j})\|^{2}\right].$
Appendix C Training Algorithm of MCBMs
Algorithm 1 Training Algorithm for MCBMs
1:Dataset $\mathcal{D}=\{\bm{x}^{(k)},\bm{y}^{(k)},\bm{c}^{(k)}\}_{k=1}^{N}$, latent norm $\lambda$, learning rate $\eta$, batch size $B$
2:Parameters $\theta$ (encoder), $\phi$ (class-head, task-heads, representation-heads)
3:Initialize parameters $\theta$, $\phi$ and representations heads:
4:for all $j=1,\dots,n$ do
5:     if $c_{j}$ is binary then
6:         $g^{z}_{j}\leftarrow\lambda\text{ if }c_{j}=1\text{ else }-\lambda$
7:     else if $c_{j}$ is multiclass then
8:         $g^{z}_{j}\leftarrow\lambda\cdot\mathrm{one\_hot}\left(c_{j}\right)$
9:     else
10:         $g^{z}_{j}\leftarrow\lambda\cdot c_{j}$
11:     end if
12:end for
13:while not converged do
14:     Sample a mini-batch $\{\bm{x}^{(k)},\bm{y}^{(k)},\bm{c}^{(k)}\}_{k=1}^{B}\sim\mathcal{D}$
15:     for all $\bm{x}^{(k)},\bm{y}^{(k)},\bm{c}^{(k)}$ in batch do
16:         Encode: Compute $\mu_{\theta}^{(k)}\leftarrow f_{\theta}\left(x^{(k)}\right)$
17:         Sample noise $\epsilon\sim\mathcal{N}(0,I)$triangleright Reparameterization trick with only one sample
18:         Reparameterize: $\bm{z}^{(k)}\leftarrow\mu_{\theta}^{(k)}+\sigma_{x}\odot\epsilon$
19:         Task prediction: $\hat{\bm{y}}^{(k)}\leftarrow g^{y}_{\phi}\left(\bm{z}^{(k)}\right)$triangleright Similar to VMs
20:         Task loss: $\mathcal{L}_{y}^{(k)}\leftarrow\left\|\bm{y}^{(k)}-\hat{\bm{y}}^{(k)}\right\|^{2}\text{ if }\bm{y}\text{ is continuous else }\text{CE}\left(\bm{y}^{(k)},\bm{\hat{y}}^{(k)}\right)$
21:         for all $j=1,\dots,n$ do
22:              Concept $j$ prediction: $\hat{c}^{(k)}_{j}\leftarrow g^{c}_{\phi,j}\left(z^{(k)}_{j}\right)$triangleright Similar to CBMs
23:              Concept $j$ loss: $\mathcal{L}_{c,j}^{(k)}\leftarrow\left\|c^{(k)}_{j}-\hat{c}^{(k)}_{j}\right\|^{2}\text{ if }c_{j}\text{ is continuous else }\text{CE}\left(c^{(k)}_{j},\hat{c}^{(k)}_{j}\right)$
24:              Representation $j$ prediction: $\hat{z}^{(k)}_{j}\leftarrow g^{z}_{j}\left(c^{(k)}_{j}\right)$triangleright Novelty in MCBMs
25:              Representation $j$ loss: $\mathcal{L}_{z,j}^{(k)}\leftarrow\left\|z^{(k)}_{j}-\hat{z}^{(k)}_{j}\right\|^{2}$
26:         end for
27:         Total loss: $\mathcal{L}^{(k)}\leftarrow\mathcal{L}_{y}^{(k)}+\beta\sum_{j=1}^{n}\mathcal{L}_{c,j}^{(k)}+\gamma\sum_{j=1}^{n}\mathcal{L}_{z,j}^{(k)}$
28:     end for
29:     Update $\theta,\phi$ using gradient descent:
$\theta\leftarrow\theta-\eta\nabla_{\theta}\left(\frac{1}{B}\sum_{k=1}^{B}\mathcal{L}^{(k)}\right),\quad\phi\leftarrow\phi-\eta\nabla_{\phi}\left(\frac{1}{B}\sum_{k=1}^{B}\mathcal{L}^{(k)}\right)$
30:end while
Appendix D Experiments Details
D.1 Hyperparameters for Section 5
Table 5: Hyperparameters for Section 5. For all datasets, the concept head $g^{c}_{\phi}$ is implemented as the identity function in CBMs, and as a multilayer perceptron (MLP) with three hidden layers in MCBMs. We emphasize that CBMs are, by design, restricted to use invertible $g^{c}_{\phi}$ to enable intervention procedures.
CIFAR-10 MPI3D Shapes3D CelebA CUB
$f_{\theta}$ architecture ResNet20 ResNet20 ResNet20 ResNet18 InceptionV3
$f_{\theta}$ pretraining None None None ImageNet ImageNet
$g^{y}_{\phi}$ hidden layers 64 64 64 64 256
low $\gamma$ 1 1 1 0.05 0.1
medium $\gamma$ 5 3 3 0.1 0.3
high $\gamma$ 10 5 5 0.3 0.5
number of epochs 50 50 50 120 250
batch size 128 128 128 128 128
optimizer SGD SGD SGD SGD SGD
learning rate $6\times 10^{-3}$ $6\times 10^{-3}$ $6\times 10^{-3}$ $2\times 10^{-2}$ $2\times 10^{-2}$
momentum 0.9 0.9 0.9 0.9 0.9
weight decay $4\times 10^{-5}$ $4\times 10^{-5}$ $4\times 10^{-5}$ $4\times 10^{-5}$ $4\times 10^{-5}$
scheduler Step Step Step Step Step
step size (epochs) 20 20 20 50 100
scheduler $\gamma$ 0.1 0.1 0.1 0.1 0.1
D.2 Datasets
CIFAR-10
The CIFAR-10 dataset consists of 60,000 color images of size 32×32, evenly distributed across 10 object classes, with 50,000 training and 10,000 test samples. Among these, 6 classes correspond to animals and 4 to vehicles. We define the target task $\bm{y}$ as a binary classification between animals and vehicles. The subset $\bm{n}_{y}$ includes one vehicle and one animal class—specifically, airplane and bird—while $\bm{c}$ comprises the remaining 8 object classes.
MPI3D
This is a synthetic dataset with controlled variation across seven generative factors: object shape, object color, object size, camera height, background color, horizontal axis, and vertical axis. In our setup, $\bm{y}$ corresponds to the object shape, $\bm{n}_{y}$ to the horizontal axis, $\bm{n}{\bar{y}}$ to the vertical axis, and $\bm{c}$ to the remaining generative factors. To ensure consistency in the mapping between concepts and the target, we filter the dataset such that any combination of elements in $\left\{\bm{c},\bm{n}_{y}\right\}$ corresponds to a unique value of $\bm{y}$. All invalid combinations are removed accordingly.
Shapes3D
This synthetic dataset consists of 3D-rendered objects placed in a room, with variation across six known generative factors: floor color, wall color, object color, scale, shape, and orientation. In our setup, $\bm{y}$ corresponds to the shape, $\bm{n}y$ includes floor color and wall color, $\bm{n}{\bar{y}}$ corresponds to orientation, and $\bm{c}$ comprises the remaining factors. We follow the same filtering strategy as in MPI3D to construct this configuration: we retain only those samples for which each combination of $\left\{\bm{c},\bm{n}_{y}\right\}$ uniquely determines $\bm{y}$, removing all invalid configurations.
CelebA
: The CelebA dataset consists of over 200,000 celebrity face images, each annotated with 40 binary facial attributes. In our setup, the task variable $\bm{y}$ corresponds to binary gender classification. For each run, eight attributes are randomly selected as the concept set $\bm{c}$, while the remaining attributes are treated as nuisances $\bm{n}_{y}$. Since most attributes exhibit some degree of correlation with gender, we set $\bm{n}_{\bar{y}}$ to the empty set.
CUB
: The Caltech-UCSD Birds (CUB) dataset contains 11,788 images of 200 bird species, annotated with part locations, bounding boxes, and 312 binary attributes. Following the approach of [9], we retain only the attributes that are present in at least 10 species (based on majority voting), resulting in a filtered set of 112 attributes. These attributes are grouped into 27 semantic clusters, where each group is defined by a common prefix in the attribute names. In our setup, the task variable $\bm{y}$ is to classify the bird species. The concept set $\bm{c}$ consists of the attributes belonging to 7 randomly selected groups (per run), while the nuisance set $\bm{n}y$ includes the attributes from the remaining 20 groups. Since most attributes exhibit some correlation with the classification task, we set $\bm{n}{\bar{y}}$ to the empty set.

### v2

@@
 This IB is implemented via a variational regularization term added to the training loss.
-As a result, MCBMs support concept-level interventions with theoretical guarantees, remain consistent with Bayesian principles, and offer greater flexibility in key design choices.
+As a result, MCBMs yield more interpretable representations, support principled concept-level interventions, and remain consistent with probability-theoretic foundations.
 1 Introduction
-Deep learning (DL) models have become the de facto standard for solving a wide range of tasks, primarily due to their ability to uncover complex patterns in data.
-However, these patterns often go beyond human comprehension, leading to DL models being perceived as black boxes—their predictions may be accurate, but they are typically difficult to interpret.
-Despite this, interpretability remains a crucial requirement in several domains, including but not limited to healthcare [1, 2, 3], finance [4, 5], and autonomous driving [6, 7].
-Most DL models operate by learning data representations—compressed versions of the input that retain essential information for tackling specific tasks of interest [8].
-Concept Bottleneck Models (CBMs) aim to define these representations based on a set human-understandable concepts [9].
-Given a task $\bm{y}$ and a set of concepts $\bm{c}=\{c_{j}\}_{j=1}^{m}$, CBMs obtain a representation $\bm{z}$ and are trained to jointly optimize the two following objectives:
-  1. (i)
-    The task $\bm{y}$ must be perfectly solved.
-  2. (ii)
-    Each concept $c_{j}$ must be predictable by using only a subset $z_{j}\in\bm{z}$.
-As a result of the second objective, CBMs are purported to offer: (i) increased interpretability of the representation space, and (ii) the ability to intervene a concrete concept $c_{j}$ by modifying $z_{j}$ and propagating these changes to the predictions.
-However, a phenomenon called information leakage has been observed in CBMs [10, 11], whereby $\bm{z}$ encodes information that is not attributable to the defined concepts.
-This issue poses two major concerns: (i) it undermines the interpretability of the learned representations, and (ii) it compromises the validity of interventions—modifying $z_{j}$ may alter not only the associated concept $c_{j}$, but also other unintended factors encoded in $z_{j}$.
-We argue that this issue stems from a fundamental limitation in the current formulation of CBMs: the absence of an explicit Information Bottleneck (IB) [12] that actively constrains $z_{j}$ to exclude information unrelated to $c_{j}$.
+Most machine learning models operate by learning data representations—compressed versions of the input that retain the essential information needed to solve a given task [1].
+However, these representations often encode information in ways that are not easily interpretable by humans.
+This lack of interpretability becomes especially problematic in sensitive domains such as healthcare [2, 3, 4], finance [5, 6], and autonomous driving [7, 8].
+To address this issue, Concept Bottleneck Models (CBMs) have been proposed, which enforce representations to be defined in terms of a set of human-understandable concepts [9].
+Formally, given a task $\bm{y}$ and a set of concepts $\bm{c}=\{c_{j}\}_{j=1}^{m}$, Vanilla Models (VMs) are trained with a single objective: their representations $\bm{z}$ should capture the information necessary to predict $\bm{y}$ accurately.
+CBMs extend this framework by adding a second objective: each concept $c_{j}$ must be recoverable from a designated component $z_{j}\in\bm{z}$.
+By enforcing this additional constraint, CBMs are claimed to offer: (i) improved interpretability of the representation space, and (ii) the ability to intervene on specific concepts by modifying $z_{j}$ and propagating these changes to the predictions.
+However, CBMs are prone to a phenomenon known as information leakage [10, 11], where the representation $\bm{z}$ encodes input information that cannot be attributed to the predefined concepts $\bm{c}$.
+We refer to this additional information as nuisances $\bm{n}$.
+Information leakage raises two main concerns: (i) it undermines interpretability, since $z_{j}$ cannot be fully explained by its corresponding concept $c_{j}$; and (ii) it compromises the validity of interventions—modifying $z_{j}$ may alter not only the associated concept $c_{j}$, but also other unintended information encoded in $z_{j}$.
+We argue that information leakage stems from a fundamental limitation in the current formulation of CBMs: the absence of an explicit Information Bottleneck (IB) [12] that actively constrains $z_{j}$ to exclude information unrelated to $c_{j}$.
 While the second objective in CBMs encourages each $z_{j}$ to retain $c_{j}$ in its entirety, it does not enforce that $z_{j}$ captures only information about $c_{j}$.
 In the worst-case scenario, $z_{j}$ could encode the entire input $\bm{x}$ and still satisfy this objective.
-We propose a solution to the aforementioned problem by introducing an IB in $z_{j}$ in such a way that it does not only preserve all the information of $c_{j}$, but it also contains information only about $c_{j}$.
-We call these models Minimal Concept Bottleneck Models (MCBMs), since $z_{j}$ is trained to be a minimal sufficient statistic of $c_{j}$ [13, 14].
-This contrasts with traditional CBMs, where $z_{j}$ is optimized to be merely a sufficient statistic of $c_{j}$.
-In other words, while CBMs simply require that $z_{j}$ preserves all information about $c_{j}$, MCBMs require that $z_{j}$ preserves only all the information about $c_{j}$.
-We implement the IB constraint via an additional loss term derived using variational approximations.
-This formulation enables faithful concept interventions without violating Bayesian principles and offers greater flexibility in key choices of the model.
-2 Related Work
-Information Leakage in Concept Bottleneck Models
-This phenomenon has been observed in CBMs trained either sequentially or jointly [10, 11], and it arises when the learned representation $\bm{z}$ encodes information that is not present in the concept set $\bm{c}$.
-Such leakage negatively impacts both the interpretability and the intervenability of the model.
-As shown in [15], this issue arises when the Markovian assumption fails—i.e., when the target $\bm{y}$ is not fully determined by the concept set $\bm{c}$, or equivalently, $p(\bm{y}|\bm{c})\neq p(\bm{y}|\bm{c},\bm{x})$, which is typically the case in real-world scenarios.
-In [16], information leakage is tackled from the perspective of out-of-distribution generalization.
-Concept Embedding Models (CEMs), introduced in [17], aim to alleviate the trade-off between accuracy and interpretability commonly observed in CBMs.
-However, we argue that this trade-off is inherently limited by the choice of concept set, and that CEMs may, in fact, be less interpretable than standard CBMs, as they yield more entropic representations—thereby amplifying information leakage.
-Other CEM-based approaches [18, 19, 20] are also prone to this issue.
-This critique is further elaborated in [21], which introduces an information-theoretic framework along with experimental benchmarks for evaluating interpretability in the representation space.
-Almost contemporaneously, [22] proposes a metric for interpretability grounded in the concept of usable mutual information [23].
-Information Bottleneck in Representation Learning
-The Information Bottleneck (IB) [12] is an information-theoretic framework that aims to balance the trade-off between preserving information about a factor variable and compressing the learned representation.
-A representation that retains all information about the factor is termed sufficient, while one that contains only information relevant to the factor—i.e., discarding all the irrelevant nuisances—is called minimal [24, 25, 26].
-This terminology originates from the theory of minimal sufficient statistics [13, 14].
-Direct computation of information-theoretic quantities is often intractable due to the complexity of the representation’s underlying distribution.
-To address this, variational approximations have been proposed to derive tractable evidence bounds that can be efficiently optimized [27, 28].
-In this work, we leverage such approximations to optimize the IB objective with respect to the concepts.
-Disentanglement
-A closely related topic to CBMs that has been extensively studied is disentanglement.
-Despite its popularity, there remains no universally agreed-upon definition.
-However, existing definitions can be broadly categorized into three groups: (i) those in [8, 29, 30, 31, 32], which define a disentangled representation as one where changes in a specific part of the representation correspond to changes in a particular factor of variation, while remaining relatively invariant to other factors; (ii) those from works like [33, 34], which view a disentangled representation as one where modifying a single factor of variation leads to changes in only one component of the representation; and (iii) more recent perspectives [35, 36, 37], which consider a representation disentangled only if it satisfies both of the above properties.
-Conventional CBMs typically yield representations that align with the second definition.
-In contrast, MCBMs encourage representations to satisfy both criteria simultaneously by introducing an Information Bottleneck for each concept, thereby aligning with the third—and currently most widely accepted—definition of disentanglement.
-Our experiments demonstrate that MCBMs produce significantly more disentangled representations.
-Table 1: Distributions considered in this work.
-$f_{\theta}$ is typically modeled by a large neural network while $g^{y}_{\phi}$, $g^{c}_{\phi}$ and $g^{z}_{\phi}$ by simpler neural networks.
-We consider $\bm{z}$ to be always continuous.
-$p_{\theta}(z|x)$ $q_{\phi}(\hat{y}|z)$ $q_{\phi}(\hat{c}_{j}|z_{j})$ $q_{\phi}(\hat{z}_{j}|c_{j})$
-Binary - $\text{Bernoulli}\left(g^{y}_{\phi}(z)\right)$ $\text{Bernoulli}\left(g^{c}_{\phi}(z_{j})\right)$
-Multiclass - $\text{Categoric}\left(g^{y}_{\phi}(z)\right)$ $\text{Categoric}\left(g^{c}_{\phi}(z_{j})\right)$
-Continuous $\mathcal{N}\left(f_{\theta}(x),\sigma^{2}_{x}I\right)$ $\mathcal{N}\left(g^{y}_{\phi}(z),\sigma_{\hat{y}}^{2}I\right)$ $\mathcal{N}\left(g^{c}_{\phi}(z_{j}),\sigma_{\hat{c}}^{2}I\right)$ $\mathcal{N}\left(g^{z}_{\phi}(c_{j}),\sigma_{\hat{z}}^{2}I\right)$
-3 An Information-Theoretic Perspective on CBMs
+Table 1: Differences between VMs, CBMs and MCBMs.
+VMs CBMs MCBMs
+Does $z_{j}$ encode all $c_{j}$?
+Does $z_{j}$ encode only $c_{j}$?
+To address this issue, we propose Minimal Concept Bottleneck Models (MCBMs), which incorporate an IB into each $z_{j}$.
+This ensures that $z_{j}$ not only retains all the information about its associated concept $c_{j}$, but also excludes any information unrelated to $c_{j}$, as summarized in Table 1.
+The name reflects that $z_{j}$ is trained to be a minimal sufficient statistic of $c_{j}$ [13, 14], in contrast to traditional CBMs where $z_{j}$ is optimized to be merely a sufficient statistic of $c_{j}$.
+As illustrated in Figure 1, this design yields disentangled representations that directly address the two shortcomings of CBMs: (i) it improves interpretability, since $z_{j}$ can be fully explained by its corresponding concept $c_{j}$; and (ii) it enables valid interventions—modifying $z_{j}$ affects only the associated concept $c_{j}$.
+In Section 2, we connect the data generative process to MCBMs through information-theoretic quantities, showing that the IB can be implemented via a variational loss.
+In Section 3, we review alternative approaches to address information leakage.
+In Section 4, we present experiments demonstrating that MCBMs enforce a true bottleneck, thereby enhancing interpretability and intervenability compared to existing alternatives.
+Finally, in Section 5, we show that the assumptions made in CBMs to enable interventions are theoretically flawed.
+Figure 1: In Vanilla Models, concepts and nuisances may be arbitrarily entangled in the representation, and a variable $z_{j}$ may capture only part of a concept (depicted as paler colors).
+In CBMs, each $z_{j}$ encodes all information about its corresponding concept $c_{j}$, but may also capture some information about nuisances (e.g., $z_{1}$) or other concepts (e.g., $z_{2}$).
+In contrast, MCBMs enforce that each representation variable $z_{j}$ encodes all—and only—the information about its corresponding concept.
+2 From Data Generative Process to MCBMs
 2.1 Data Generative Process
@@
 We refer to models trained to address this problem as Vanilla Models (VMs).
-These models typically operate by first extracting an intermediate representation $\bm{z}$ from the input $\bm{x}$ via an encoder $p_{\theta}(\bm{z}\mid\bm{x})$.
+These models typically operate by first extracting an intermediate representation $\bm{z}\in\mathcal{Z}$ from the input $\bm{x}$ via an encoder $p_{\theta}(\bm{z}\mid\bm{x})$.
 Subsequently, a prediction $\hat{y}\in\mathcal{Y}$ is produced from $\bm{z}$ using a task head $q_{\phi}(\hat{\bm{y}}\mid\bm{z})$.
@@
 The choice of task head $q_{\phi}(\hat{\bm{y}}\mid\bm{z})$ depends on the structure of the output space $\mathcal{Y}$, also detailed in Table 2.
-The objective in Equation (1) corresponds to minimizing the cross-entropy loss when $\bm{y}$ is binary or multiclass, and to minimizing the mean squared error between $\bm{y}$ and $g^{y}_{\phi}(z)$ when $\bm{y}$ is continuous.
+The objective in Equation 1 corresponds to minimizing the cross-entropy loss when $\bm{y}$ is binary or multiclass, and to minimizing the mean squared error between $\bm{y}$ and $g^{y}_{\phi}(z)$ when $\bm{y}$ is continuous.
 2.3 Concept Bottleneck Models
@@
 As illustrated in Figure 2(c), CBMs extend Vanilla Models by incorporating a concept head $q(\hat{c}_{j}\mid z_{j})$, depicted with blue arrows.
-These models jointly optimize the objectives in Equations (1) and (2).
+These models jointly optimize the objectives in Equations 1 and 2.
 As detailed in Table 2, the form of the concept head $q_{\phi}(\hat{c}_{j}\mid z_{j})$ depends on the nature of the concept space $\mathcal{C}$.
-The objective in Equation (2) corresponds to minimizing the cross-entropy loss when $c_{j}$ is binary or multiclass, and the mean squared error between $c_{j}$ and $g_{\phi}^{c}(z_{j})$ when $c_{j}$ is continuous.
-(a) Data Generative Process
-(b) Vanilla Model (VM)
-(c) Concept Bottleneck Model (CBM)
-(d) Minimal Concept Bottleneck Model (MCBM)
-Figure 1: Graphical models of the different systems described.
-We consider two concepts and two-dimensional representations.
-We provide in Appendix A the analogous figure for $m$ concepts and $m$-dimensional representations.
-Inputs $\bm{x}$ are defined by some concepts $\{c_{j}\}_{j=1}^{m}$ and nuisances $\bm{n}$; and targets $\bm{y}$ are defined by $\bm{x}$ (gray arrows).
-Vanilla models obtain the representations $\{z_{j}\}_{j=1}^{m}$ from $\bm{x}$ through the encoder $p_{\theta}(\bm{z}|\bm{x})$ (black arrows) and solve the task $\hat{\bm{y}}$ sequentially through the task head $q_{\phi}(\bm{y}|\bm{x})$ (green arrows).
-Concept Bottleneck Models try to make a prediction $\hat{c}_{j}$ of each concept $c_{j}$ from one representation $z_{j}$ through the concept head $q(\hat{c}_{j}|z_{j})$ (blue arrows).
-Minimal CBMs (ours) try to make a prediction $z_{j}$ of each representation $z_{j}$ from one concept $c_{j}$ through the representation head $q(z_{j}|c_{j})$ (red arrows).
-4 On the Lack of a Bottleneck in CBMs
-As explained in Section 1, one of the main advantages commonly attributed to CBMs over vanilla models is their ability to perform concept-level interventions and propagate the effects to the model’s predictions.
-For instance, consider a scenario in which we aim to estimate $p(\hat{\bm{y}}\mid c_{j}=\alpha,\bm{x})$—that is, the model’s predicted output $\hat{\bm{y}}$ had the concept $c_{j}$ been set to the value $\alpha$.
-In CBMs, this intervention is typically performed through the latent representation $z_{j}$, according to:
+The objective in Equation 2 corresponds to minimizing the cross-entropy loss when $c_{j}$ is binary or multiclass, and the mean squared error between $c_{j}$ and $g_{\phi}^{c}(z_{j})$ when $c_{j}$ is continuous.
+How are Interventions Performed in CBMs?
+As explained in Section 1, a key advantage often attributed to CBMs is their ability to support concept-level interventions.
+Suppose we aim to estimate $p(\hat{\bm{y}}\mid c_{j}=\alpha,\bm{x})$.
+In CBMs, this intervention is performed through the latent representation $z_{j}$:
 […]
-However, in the original formulation of CBMs (described in Section 3.3), the conditional distribution $p(z_{j}\mid c_{j})$ is not defined—there is no directed path from $c_{j}$ to $z_{j}$ in Figure 1(c).
+However, the conditional distribution $p(z_{j}\mid c_{j})$ is not defined—there is no directed path from $c_{j}$ to $z_{j}$ in Figure 2(c).
 Intuitively, because $z_{j}$ may encode information about $\bm{x}$ beyond $c_{j}$, it cannot be fully determined by $c_{j}$ alone.
-This leads to a key question: how can interventions be performed in CBMs if $p(z_{j}\mid c_{j})$ is unknown?
-In what follows, we address this question and argue that current practices often violate basic Bayesian principles, leading to inconsistencies and limiting the architectural flexibility of the model.
-To resolve this, we introduce Minimal CBMs (MCBMs), a theoretically grounded framework that addresses these issues and enables principled interventions.
-4.1 Intervention Mechanisms in Concept Bottleneck Models
-As previously discussed, the conditional distribution $p(z_{j}\mid c_{j})$ is not explicitly defined in standard CBMs.
-To enable a mechanism for performing interventions, two constraints are typically imposed:
+This raises a key question: how can interventions be performed in CBMs if $p(z_{j}\mid c_{j})$ is unknown?
+To make interventions feasible, CBMs typically impose two constraints:
   1. (i)
-    Concepts $c_{j}$, for $j=1,2,\dots,m$, are assumed to be binary. If a concept is originally multiclass with $k$ categories, it is converted into $k$ binary concepts (One-vs-Rest [38]).
+    Concepts $c_{j}\in\bm{c}$, are assumed to be binary. If a concept is originally multiclass with $k$ categories, it is converted into $k$ binary concepts (One-vs-Rest [15]).
   2. (ii)
     The concept head is defined as $q_{\phi}(c_{j}\mid z_{j})=\sigma(z_{j})$, where $\sigma$ denotes the sigmoid function.
-Given that $\sigma$ is an invertible function, this setup permits defining $p(z_{j}\mid c_{j})=\sigma^{-1}(c_{j})$.
-However, since $\sigma^{-1}(1)=-\sigma^{-1}(0)=\infty$, this inversion becomes ill-defined at the binary extremes.
-To address this, in practice, $p(z_{j}\mid c_{j}=0)$ and $p(z_{j}\mid c_{j}=1)$ are approximated as the 5th and 95th percentiles of the empirical distribution of $z_{j}$, respectively [9].
-Nevertheless, this practice introduces two issues:
-  1. 1.
-    One-vs-Rest strategies exhibit the following limitations: (i) individual binary classifiers tend to be biased toward the negative class, (ii) multiple classifiers may simultaneously assign a "positive" label to the same input, and (iii) the predicted probabilities across classifiers are typically uncalibrated [39].
-These limitations are further illustrated in Section 5.3.
-  2. 2.
-    The intervention procedure based on the sigmoid inverse does not adhere to Bayes’ rule, i.e., $p(z_{j}\mid c_{j})=\frac{q_{\phi}(c_{j}\mid z_{j})p(z_{j})}{p(c_{j})}$.
-While $p(c_{j})$ can be approximated using its empirical distribution, the marginal $p(z_{j})$ remains unknown and is challenging to estimate.
+Since $\sigma$ is invertible, this setup permits defining $p(z_{j}\mid c_{j})\approx\sigma^{-1}(c_{j})$.
+However, this is ill-defined at the binary extremes, as $\sigma^{-1}(1)=-\sigma^{-1}(0)=\infty$.
+To address this, in practice, $p(z_{j}\mid c_{j}=0)$ and $p(z_{j}\mid c_{j}=1)$ are set as the 5th and 95th percentiles of the empirical distribution of $z_{j}$, respectively [9].
+This workaround, however, introduces two crucial issues discussed in Section 5.
+Table 2: Distributions considered in this work.
+$f_{\theta}$ is typically modeled by a large neural network while $g^{y}_{\phi}$, $g^{c}_{\phi}$ and $g^{z}_{\phi}$ by simpler neural networks.
+We consider $\bm{z}$ to be always continuous.
+$p_{\theta}(z|x)$ $q_{\phi}(\hat{y}|z)$ $q_{\phi}(\hat{c}_{j}|z_{j})$ $q_{\phi}(\hat{z}_{j}|c_{j})$
+Binary - $\text{Bernoulli}\left(g^{y}_{\phi}(z)\right)$ $\text{Bernoulli}\left(g^{c}_{\phi}(z_{j})\right)$ -
+Multiclass - $\text{Categoric}\left(g^{y}_{\phi}(z)\right)$ $\text{Categoric}\left(g^{c}_{\phi}(z_{j})\right)$ -
+Continuous $\mathcal{N}\left(f_{\theta}(x),\sigma^{2}_{x}I\right)$ $\mathcal{N}\left(g^{y}_{\phi}(z),\sigma_{\hat{y}}^{2}I\right)$ $\mathcal{N}\left(g^{c}_{\phi}(z_{j}),\sigma_{\hat{c}}^{2}I\right)$ $\mathcal{N}\left(g^{z}_{\phi}(c_{j}),\sigma_{\hat{z}}^{2}I\right)$
 2.4 Minimal Concept Bottleneck Models
-As previously discussed, CBMs lack an explicit mechanism for enforcing a bottleneck, which often undermines their primary advantages.
+As discussed in Section 1, CBMs lack an explicit mechanism for enforcing a bottleneck, which often undermines their intended advantages.
 To address this limitation, we introduce Minimal Concept Bottleneck Models (MCBMs), which explicitly impose an Information Bottleneck.
-This ensures that each representation $z_{j}$ not only retains all information about the associated concept $c_{j}$, but also contains information only about $c_{j}$.
+This ensures that each $z_{j}$ retains all information about its associated concept $c_{j}$, while excluding information unrelated to $c_{j}$.
 In other words, $z_{j}$ becomes a minimal sufficient representation of $c_{j}$.
-To achieve this, we introduce a representation head $q_{\phi}(\hat{z}j\mid c_{j})$ that predicts $z_{j}$ from $c_{j}$.
+To achieve this, we introduce a representation head $q_{\phi}(\hat{z}_{j}\mid c_{j})$ that predicts $z_{j}$ from $c_{j}$.
 This encourages $z_{j}$ to discard information unrelated to $c_{j}$, as doing so improves the predictive accuracy of $q_{\phi}(\hat{z}_{j}\mid c_{j})$.
@@
 Figure 2(d) illustrates how MCBMs extend CBMs by introducing the representation head $q(\hat{z}_{j}\mid c_{j})$, depicted with red arrows.
-These models are trained to jointly optimize the objectives given in Equations (1), (2), and (4).
-Although the KL Divergence in Equation (4) does not admit a closed-form solution in general, we show in Appendix B.4 that, under the distributional assumptions listed in Table 1, it reduces to the mean squared error between $f_{\theta}(x)$ and $g^{z}_{\phi}(c_{j})$.
-4.3 Practical Considerations for Optimizing MCBMs
-To optimize MCBMs, we combine the three previously introduced objectives, incorporating two key considerations.
-First, we approximate the expectations over $p(\bm{x},\bm{y})$ and $p(\bm{x},c_{j})$ using the empirical data distribution described in Section 3.1, replacing integrals with summations over the dataset.
-Second, to enable gradient-based optimization through the stochastic encoder, we apply the reparameterization trick [40], allowing for low-variance gradient estimates during backpropagation.
-$\mathbb{E}_{p_{\theta}(\bm{z}|\bm{x})}\left[\log q_{\phi}(\bm{y}|\bm{z})\right]\approx\sum_{i}\log q_{\phi}\left(\bm{y}|f^{\prime}_{\theta}\left(x,\epsilon^{(i)}\right)\right)$ and $E_{p_{\theta}(z_{j}|\bm{x})}\left[\log q_{\phi}(\hat{c}_{j}|z_{j})\right]\approx\sum_{i}\log q_{\phi}\left(\hat{c}_{j}\big{|}f^{\prime}_{\theta,j}\left(\bm{x},\epsilon^{(i)}\right)\right)$, where $f^{\prime}_{\theta}(\bm{x},\epsilon)=f_{\theta}(\bm{x})+\sigma^{2}_{x}I\epsilon$ (due to the choice of $p_{\theta}(\bm{z}|\bm{x})$ in Table 1), $\epsilon\sim\mathcal{N}\left(0,I\right)$ and $f^{\prime}_{\theta,j}\left(x,\epsilon\right)$ corresponds to the element $j$ of $f^{\prime}_{\theta}\left(\bm{x},\epsilon\right)$.
-Combining the considerations above yields the final objective for training MCBMs, as shown in Equation (5), where $\beta$ and $\gamma$ are hyperparameters.
-The first term corresponds to the objective used in Vanilla Models, the second term is introduced in CBMs, and the third is specific to MCBMs.
-s.
-Detailed training algorithms for the various cases listed in Table 1 are provided in Appendix C.
-$\begin{split}\max_{\theta,\phi}\sum_{k=1}^{N}\sum_{i}\log q_{\phi}\left(\hat{\bm{y}}\big{|}f^{\prime}_{\theta}\left(x^{(k)},\epsilon^{(i)}\right)\right)+&\beta\sum_{j=1}^{n}\log q_{\phi}\left(\hat{c}_{j}\big{|}f^{\prime}_{\theta,j}\left(\bm{x}^{(k)},\epsilon^{(i)}\right)\right)\\ &-\gamma\sum_{j=1}^{n}D_{KL}\left(p_{\theta}\left(z_{j}|\bm{x}^{(k)}\right)\big{|}\big{|}\ q_{\phi}\left(\hat{z}_{j}|c_{j}^{(k)}\right)\right)\end{split}$ (5)
-4.4 Intervention Mechanisms in Minimal Concept Bottleneck Models
-As discussed at the beginning of this section, intervention procedures require access to $p(z_{j}\mid c_{j})$.
-As outlined in Section 4.1, this dependence anchors the intervention process to specific architectural design choices in CBMs.
-Moreover, such procedures violate Bayesian principles, as the prior distribution $p(z_{j})$ is typically unknown.
-In MCBMs, since $z_{j}$ is explicitly designed to contain information only about $c_{j}$, modifying $z_{j}$ corresponds to intervening solely on $c_{j}$.
-This is reflected in the graphical model illustrated in Figure 1(d), where MCBMs introduce a directed path from $c_{j}$ to $z_{j}$ through the intermediate variable $\hat{z}_{j}$.
+These models are trained to jointly optimize the objectives given in Equations 1, 2, and 4.
+Although the KL Divergence in Equation 4 does not admit a closed-form solution in general, we show in Appendix B.4 that, under the distributional assumptions listed in Table 2, it reduces to the mean squared error between $f_{\theta}(x)$ and $g^{z}_{\phi}(c_{j})$.
+How are Interventions Performed in MCBMs?
+In contrast to CBMs, MCBMs explicitly constrain $z_{j}$ to contain only information about $c_{j}$.
+As a result, modifying $z_{j}$ corresponds to intervening solely on $c_{j}$.
+This is reflected in Figure 2(d), where MCBMs introduce a directed path from $c_{j}$ to $z_{j}$ through the intermediate variable $\hat{z}_{j}$.
 This structure permits the computation of:
-[…]
-which simplifies to $p(z_{j}\mid c_{j})=q_{\phi}(z_{j}\mid c_{j})$ when $p(z_{j}\mid\hat{z}_{j})=\delta(z_{j}-\hat{z}_{j})$, which holds when the objective in Equation (4) is optimized to convergence—or, equivalently, when $z_{j}$ contains information exclusively about $c_{j}$.
-As shown in Table 1, we define the encoder distribution as $q_{\phi}(z_{j}\mid c_{j})=\mathcal{N}(g^{z}_{\phi}(c_{j}),\sigma_{\hat{z}}^{2}I)$.
-The mean function $g^{z}_{\phi}(c_{j})$ is chosen according to the following rules:
+$p(z_{j}|c_{j})=\int p(z_{j}|\hat{z}_{j})q_{\phi}(\hat{z}_{j}|c_{j})\,d\hat{z}_{j}$ (5)
+which simplifies to $p(z_{j}\mid c_{j})=q_{\phi}(z_{j}\mid c_{j})$ when $p(z_{j}\mid\hat{z}_{j})=\delta(z_{j}-\hat{z}_{j})$, i.e., once the objective in Equation 4 is optimized and $z_{j}$ encodes exclusively $c_{j}$.
+As shown in Table 2, we define the encoder distribution as $q_{\phi}(z_{j}\mid c_{j})=\mathcal{N}(g^{z}_{\phi}(c_{j}),\sigma_{\hat{z}}^{2}I)$, where mean function $g^{z}_{\phi}(c_{j})$ is chosen according to the following rules:
   1. (i)
@@
     For categorical concepts, $g^{z}_{\phi}(c_{j})=\lambda\cdot\mathrm{one\_hot}(c_{j})$.
-This mirrors the Prototypical Learning [41] approach where class-dependent prototypes $\{g^{z}_{\phi}(c_{j})\}_{j=1}^{m}$ are fixed, as adopted in [42].
+This mirrors the Prototypical Learning [16] approach where class-dependent prototypes $\{g^{z}_{\phi}(c_{j})\}_{j=1}^{m}$ are fixed.
   3. (iii)
     For continuous concepts, $g^{z}_{\phi}(c_{j})=\lambda\cdot c_{j}$.
-Here, $\lambda$ is a scaling constant that controls the norm of the latent representation and is fixed to $\lambda=3$ across all experiments.
+Here, $\lambda$ is a scaling constant that controls the norm of the latent representation, fixed to $\lambda=3$ in all experiments.
 Regarding the variance term, we set: (i) $\sigma_{x}=0$in the case of CBMs to obtain a deterministic encoder in line with their original formulation, and (ii) $\sigma_{x}=\sigma_{\hat{z}}=1$for MCBMs.
+Practical Considerations for Optimizing MCBMs
+To optimize MCBMs, we combine the three previously introduced objectives, incorporating two key considerations.
+First, we approximate the expectations over $p(\bm{x},\bm{y})$ and $p(\bm{x},c_{j})$ using the empirical data distribution, replacing integrals with summations over the dataset.
+Second, to enable gradient-based optimization through the stochastic encoder, we apply the reparameterization trick [17]: $\mathbb{E}_{p_{\theta}(\bm{z}|\bm{x})}\left[\log q_{\phi}(\bm{y}|\bm{z})\right]\approx\sum_{i}\log q_{\phi}\left(\bm{y}|f^{\prime}_{\theta}\left(x,\epsilon^{(i)}\right)\right)$ and $E_{p_{\theta}(z_{j}|\bm{x})}\left[\log q_{\phi}(\hat{c}_{j}|z_{j})\right]\approx\sum_{i}\log q_{\phi}\left(\hat{c}_{j}\big|f^{\prime}_{\theta,j}\left(\bm{x},\epsilon^{(i)}\right)\right)$, where $f^{\prime}_{\theta}(\bm{x},\epsilon)=f_{\theta}(\bm{x})+\sigma^{2}_{x}I\epsilon$ (due to the choice of $p_{\theta}(\bm{z}|\bm{x})$ in Table 2), $\epsilon\sim\mathcal{N}\left(0,I\right)$ and $f^{\prime}_{\theta,j}\left(x,\epsilon\right)$ corresponds to the element $j$ of $f^{\prime}_{\theta}\left(\bm{x},\epsilon\right)$.
+Combining the considerations above yields the final objective for training MCBMs, as shown in Equation 6, where $\beta$ and $\gamma$ are hyperparameters.
+The first term corresponds to the objective used in Vanilla Models, the second term is introduced in CBMs, and the third is specific to MCBMs.
+Detailed training algorithms for the various cases listed in Table 2 are provided in Appendix C.
+$\begin{split}\max_{\theta,\phi}\sum_{k=1}^{N}\sum_{i}\log q_{\phi}\left(\hat{\bm{y}}\big|f^{\prime}_{\theta}\left(x^{(k)},\epsilon^{(i)}\right)\right)+&\beta\sum_{j=1}^{n}\log q_{\phi}\left(\hat{c}_{j}\big|f^{\prime}_{\theta,j}\left(\bm{x}^{(k)},\epsilon^{(i)}\right)\right)\\ &-\gamma\sum_{j=1}^{n}D_{KL}\left(p_{\theta}\left(z_{j}|\bm{x}^{(k)}\right)\big|\big|\ q_{\phi}\left(\hat{z}_{j}|c_{j}^{(k)}\right)\right)\end{split}$ (6)
+(a) Data Generative Process
+(b) Vanilla Model (VM)
+(c) Concept Bottleneck Model (CBM)
+(d) Minimal Concept Bottleneck Model (MCBM)
+Figure 2: Graphical models of the different systems described for two concepts and two-dimensional representations.
+We provide in Appendix A the analogous figure for $m$ concepts and $m$-dimensional representations.
+Inputs $\bm{x}$ are defined by some concepts $\{c_{j}\}_{j=1}^{m}$ and nuisances $\bm{n}$; and targets $\bm{y}$ are defined by $\bm{x}$ (gray arrows).
+Vanilla models obtain the representations $\{z_{j}\}_{j=1}^{m}$ from $\bm{x}$ through the encoder $\color[rgb]{0,1,1}p_{\theta}(\bm{z}|\bm{x})$ (cyan arrows) and solve the task $\hat{\bm{y}}$ sequentially through the task head $\color[rgb]{0,1,0}q_{\phi}(\hat{\bm{y}}|\bm{z})$ (green arrows).
+Concept Bottleneck Models make a prediction $\hat{c}_{j}$ of each concept $c_{j}$ from one representation $z_{j}$ through the concept head $\color[rgb]{0,0.49609375,1}q(\hat{c}_{j}|z_{j})$ (blue arrows).
+Minimal CBMs make a prediction $z_{j}$ of each representation $z_{j}$ from one concept $c_{j}$ through the representation head $\color[rgb]{1,0,0}q(z_{j}|c_{j})$ (red arrows).
+3 Related Work
+Information Leakage in Concept Bottleneck Models
+Information leakage occurs when the learned representation $\bm{z}$ encodes information outside the concept set $\bm{c}$, reducing both interpretability and intervenability [10, 11].
+This arises when the Markovian assumption fails—i.e., when the target $\bm{y}$ is not fully determined by the concept set $\bm{c}$, or equivalently, $p(\bm{y}|\bm{c})\neq p(\bm{y}|\bm{c},\bm{x})$, which is typically the case in real-world scenarios [18].
+Concept Embedding Models (CEMs) [19] were proposed to mitigate the accuracy–interpretability trade-off in CBMs.
+However, this trade-off is fundamentally limited by the chosen concept set, and CEMs may even be less interpretable than standard CBMs, as their more entropic representations tend to amplify information leakage.
+Recently, this critique has been further formalized: Parisini et al.
+[20] proposed information-theoretic benchmarks for assessing interpretability and Makonnen et al.
+[21] introduced new metrics grounded in usable mutual information [22].
+Havasi et al.
+[18] also introduced Hard Concept Bottleneck Models (HCBMs), which predict $\bm{y}$ from binarized concept predictions $\hat{c}_{j}$ rather than from $\bm{z}$ (see Appendix D), thereby imposing an ad-hoc Information Bottleneck.
+Extensions such as Autoregressive CBMs and Stochastic CBMs [18, 23] incorporate dependencies between concepts.
+Information Bottleneck in Representation Learning
+The Information Bottleneck (IB) [12] provides a principled way to balance preserving information about a factor with compressing the representation.
+In this framework, a representation is sufficient if it retains all the information about the variable of interest, and minimal if it contains only that information [24, 25, 26].
+Computation of these quantities is intractable, so variational methods have been developed to derive tractable evidence bounds [27, 28].
 4 Experiments
@@
 In contrast, we show that MCBMs successfully enforce a bottleneck, thereby mitigating the issues that arise in its absence.
-Additional implementation details, including encoder architectures and training hyperparameters for each experiment, are provided in Appendix D.
+Further implementation details—including encoder architectures and training hyperparameters for each experiment—are provided in Appendix E and at https://github.com/antonioalmudevar/minimal_cbm.
 4.1 Do CBMs and MCBMs Leak Information?
@@
 None of them should affect the representation $\bm{z}$ since it must be fully described by the concepts $\bm{c}$.
-In some works, it is assumed that the task $\bm{y}$ is fully defined by the concepts $\bm{c}$.
-However, it is unrealistic to have a finite set of concepts that we understand as humans and that totally describe an arbitrarily complex task.
-In general, some of the nuisances $\bm{n}$ that describe the input $\bm{x}$ will also influence $\bm{y}$.
-Specifically, we decompose the nuisances as $\bm{n}=\{\bm{n}_{y},\bm{n}_{\bar{y}}\}$, where $\bm{n}_{y}$ captures nuisances that, together with the concepts, describe $\bm{y}$, while $\bm{n}_{\bar{y}}$ comprises those that are independent of $\bm{y}$ (see Figure 2(a)).
-As shown through this text, CBMs have no incentive to remove information that is not related to the concepts.
-Additionally, since the objective of models in general is to solve $\bm{y}$, they have incentives to preserve not only $\bm{c}$ but also $\bm{n}_{y}$ in their representations.
-Although this may lead to improved performance in solving the task, it comes at the expense of interpretability and the ability to intervene in the system.
-For example if $z_{j}\in\bm{z}$ were intended to represent the concept $c_{j}\in\bm{c}$, then one could believe that an intervention is made solely on $c_{j}$ by modifying $z_{j}$.
-However $z_{j}$ could also contain information on $\bm{n}_{y}$ (or even $\bm{n}_{\bar{y}}$, as we see later), so we would be intervening also the nuisances by modifying $z_{j}$, thereby invalidating any causal conclusion drawn.
-By introducing the Information Bottleneck in equation (4), we force the model to remove the nuisances from $z_{j}$, thus restoring the validity of any causal analysis about $c_{j}$.
-We note that this will necessarily negatively impact the performance in solving the task $\bm{y}$ by the nature of the problem.
-If $\bm{c}$ is incomplete and information from $\bm{n}$ is necessary to solve $\bm{y}$ (i.e., $\bm{n}_{\bar{y}}\neq\emptyset$), the performance will decrease if $\bm{z}$ does not contain information about $\bm{n}$.
-We analyze next i $\bm{n}_{y}$ and $\bm{n}_{\bar{y}}$ are present in the representation for different versions of CBMs and datasets.
+Some works assume that the task $\bm{y}$ is fully defined by the concepts $\bm{c}$.
+However, it is unrealistic to expect a finite set of human-understandable concepts to completely describe an arbitrarily complex task.
+In practice, certain nuisances $\bm{n}$ that describe the input $\bm{x}$ also influence $\bm{y}$.
+Specifically, we decompose the nuisances as $\bm{n}=\{\bm{n}_{y},\bm{n}_{\bar{y}}\}$, where $\bm{n}_{y}$ captures nuisances that, together with the concepts, describe $\bm{y}$, while $\bm{n}_{\bar{y}}$ comprises those independent of $\bm{y}$ (see Figure 3(a)).
+As discussed throughout this work, CBMs provide no incentive to remove information unrelated to the concepts.
+Moreover, since the objective of most models is to solve $\bm{y}$, they are incentivized to preserve not only $\bm{c}$ but also $\bm{n}_{y}$.
+While this may improve task performance, it comes at the expense of interpretability and valid interventions.
+For instance, if $z_{j}\in\bm{z}$ is intended to represent the concept $c_{j}\in\bm{c}$, , one might assume that modifying $z_{j}$ corresponds solely to intervening on $c_{j}$.
+However, if $z_{j}$ also encodes information about $\bm{n}_{y}$ (or even $\bm{n}_{\bar{y}}$ as we show later), then modifying $z_{j}$ also affects the nuisances, invalidating any causal conclusions.
+By introducing the Information Bottleneck in Equation 4, we explicitly constrain the model to remove nuisances from $z_{j}$, thereby restoring the validity of causal analyses involving $c_{j}$.
+We note, however, that this necessarily reduces task performance: if $\bm{c}$ is incomplete and solving $\bm{y}$ requires information from $\bm{n}$ (i.e., $\bm{n}_{y}\neq\emptyset$), excluding $\bm{n}$ from $\bm{z}$ will lower predictive performance.
+We next examine whether $\bm{n}_{y}$ and $\bm{n}_{\bar{y}}$ are present in the representations across different CBM variants and datasets.
+For this purpose, we define the following task–concept configurations: (i) MPI3D[29], where $\bm{y}$ is the object shape, $\bm{n}_{y}$ the horizontal axis, $\bm{n}_{\bar{y}}$ the vertical axis, and $\bm{c}$ the remaining generative factors; (ii) Shapes3D[30], where $\bm{y}$ is the shape, $\bm{n}_{y}$ the floor color and wall color, $\bm{n}_{\bar{y}}$ the orientation, and $\bm{c}$ the remaining factors; (iii) CIFAR-10[31], where $\bm{y}$ is the standard classification task, $\bm{c}$ consists of 64 of the 143 attributes extracted by Oikarinen et al.
+[32] using GPT-3 [33], and $\bm{n}_{y}$ the remaining attributes, since all nuisances are correlated with $\bm{y}$; and (iv) CUB[34], where $\bm{y}$ is the bird species, $\bm{c}$ includes concepts from twelve randomly selected attribute groups, and $\bm{n}_{y}$ the attributes from the remaining 20 groups.
+While MCBMs natively support multiclass concepts, the other baselines in our study are limited to binary concepts.
+To ensure fairness, factors with $k$ classes are therefore represented as $k$ binary concepts.
 Table 3: Average value of URR for task-related nuisances $\bm{n}_{y}$.
-CIFAR-10 MPI3D Shapes3D CelebA CUB
-Vanilla $33.5\pm 4.3$ $35.0\pm 1.9$ $45.5\pm 6.4$ $10.7\pm 3.5$ $3.8\pm 1.0$
-CBM $34.7\pm 0.2$ $28.1\pm 0.5$ $8.1\pm 2.9$ $11.5\pm 3.5$ $3.8\pm 0.8$
-CEM $35.1\pm 0.1$ $43.2\pm 5.2$ $15.8\pm 3.9$ $16.8\pm 3.3$ $3.9\pm 1.1$
-MCBM (low $\gamma$) $24.8\pm 0.6$ $10.7\pm 0.1$ $2.5\pm 0.1$ $4.3\pm 0.0$ $3.4\pm 0.9$
-MCBM (medium $\gamma$) $19.1\pm 1.0$ $6.7\pm 0.2$ $0.2\pm 0.3$ $\bm{7.9\pm 2.9}$ $2.8\pm 0.8$
-MCBM (high $\gamma$) $\bm{11.0\pm 0.5}$ $\bm{0.0\pm 0.0}$ $\bm{0.0\pm 0.0}$ $8.0\pm 2.4$ $\bm{2.4\pm 1.0}$
+MPI3D Shapes3D CIFAR-10 CUB
+Vanilla $35.0\pm 1.9$ $45.5\pm 6.4$ $19.8\pm 0.7$ $3.8\pm 1.0$
+CBM [9] $28.1\pm 0.5$ $18.1\pm 2.9$ $18.5\pm 0.7$ $3.8\pm 0.8$
+CEM [19] $43.2\pm 5.2$ $15.8\pm 3.9$ $27.2\pm 0.8$ $3.9\pm 1.1$
+ARCBM [18] $28.2\pm 1.7$ $18.4\pm 2.1$ $18.2\pm 0.6$ $3.9\pm 0.9$
+SCBM [23] $24.3\pm 0.4$ $21.8\pm 1.4$ $18.3\pm 0.7$ $3.6\pm 0.8$
+MCBM (low $\gamma$) $10.7\pm 0.1$ $2.5\pm 0.1$ $18.0\pm 0.5$ $3.4\pm 0.9$
+MCBM (medium $\gamma$) $6.7\pm 0.2$ $0.2\pm 0.3$ $18.1\pm 0.5$ $2.8\pm 0.8$
+MCBM (high $\gamma$) $\bm{0.0\pm 0.0}$ $\bm{0.0\pm 0.0}$ $\bm{17.6\pm 0.5}$ $\bm{2.4\pm 1.0}$
 Task-related information leakage
-(a) CBM ($\beta=1$)
-(b) MCBM ($\beta=1$, $\gamma=5$)
-Figure 3: Average prediction $\hat{y}$ in animal (0) vs. vehicle (1) classification for the different classes when excluding airplane and bird from the concept set.
-(c) CBM ($\beta=1$)
-(d) MCBM ($\beta=1$, $\gamma=5$)
-Figure 4: Average prediction $\hat{y}$ for animal (purple) vs. vehicle (yellow) when intervening on $\bm{z}$ to match the x-axis class instead of the true class (y-axis).
-We analyze whether $\bm{n}_{y}$ is encoded in $\bm{z}$—that is, whether the dependency illustrated by the orange arrow in Figure 2(b) is present—and examine its consequences.
-To illustrate this, we design a simple experiment based on the CIFAR-10 dataset [43].
-Since all classes in CIFAR-10 correspond to either animals or vehicles, we define the target $\bm{y}$ as the binary task of distinguishing between these two categories.
-The concepts $\bm{c}$ are defined using all classes except airplane and bird.
-Under a proper bottleneck, the model should be unable to accurately predict $\bm{y}$ for these excluded classes, as their information is not captured by any concept.
-Figure 4 shows the mean predictions $\hat{y}$ for each class, comparing a CBM and an MCBM.
-The CBM confidently predicts $\bm{y}$ even for the excluded classes, indicating a substantial degree of information leakage.
-In contrast, the MCBM significantly mitigates this issue, displaying low confidence for excluded classes while maintaining high confidence for those included in the concept set.
-To assess the impact on interventions, Figure 4 shows model predictions after simulating an intervention corresponding to each concept.
-In the CBM, predictions following interventions exhibit reduced confidence for certain classes—an effect attributed to residual nuisance information in the representation.
-This effect is notably reduced in the MCBM.
-To systematically analyze this phenomenon, we consider several datasets and define distinct task–concept configurations for each.
-Specifically: (i) In CIFAR-10 [43], we adopt the described scenario: $\bm{y}$ corresponds to a binary classification between animals and vehicles, $\bm{n}_{y}$ includes airplane and bird, and $\bm{c}$ represents the remaining classes.
-(ii) In MPI3D [44], $\bm{y}$ corresponds to the object shape, $\bm{n}_{y}$ to the horizontal axis, $\bm{n}_{\bar{y}}$ to the vertical axis, and $\bm{c}$ to the remaining generative factors.
-(iii) In Shapes3D [30], $\bm{y}$ is the shape, $\bm{n}_{y}$ includes the floor color and wall color, $\bm{n}_{\bar{y}}$ is the orientation, and $\bm{c}$ includes the remaining factors.
-(iv) In CelebA [45], $\bm{y}$ is gender, $\bm{c}$ consists of eight randomly selected attributes, and $\bm{n}_{y}$ includes the remaining factors, all of which are correlated with $\bm{y}$.
-(v) In CUB [46], $\bm{y}$ corresponds to the bird species, $\bm{c}$ includes the concepts from seven randomly selected attribute groups, and $\bm{n}_{y}$ comprises the attributes from the remaining 20 groups.
-Further details for each setup are provided in Appendix D.
-We recall that, while our MCBMs natively support multiclass concepts, the other models evaluated in this work are limited to binary concepts.
-Therefore, to ensure fair comparisons, factors with $m$ classes are represented as $m$ separate binary concepts.
-To assess the presence of a nuisance factor $n_{j}\in\bm{n}_{y}$ in $\bm{z}$, we aim to estimate $I(N_{j};Z\mid C)$—that is, the amount of additional information that $\bm{z}$ provides about $n_{j}$ beyond what is already captured by $\bm{c}$.
-Since this quantity is intractable in practice, we approximate it as $\hat{I}(N_{j};Z|C)=\hat{H}(N_{j}|C)-\hat{H}(N_{j}|C,Z)$, where $\hat{H}(N_{j}\mid C)=-\sum_{k=1}^{N}\log h_{\psi}^{c}(\bm{c}^{(k)})$ and $\hat{H}(N_{j}\mid C,Z)=-\sum_{k=1}^{N}\log h_{\psi}^{cz}(\bm{c}^{(k)},\bm{z}^{(k)})$.
+To measure the presence of a nuisance factor $n_{j}\in\bm{n}_{y}$ in $\bm{z}$, we estimate $I(N_{j};Z\mid C)$—the information about $n_{j}$ contained in $\bm{z}$ beyond what is explained by $\bm{c}$.
+Since this quantity is intractable, we approximate it as $\hat{I}(N_{j};Z|C)=\hat{H}(N_{j}|C)-\hat{H}(N_{j}|C,Z)$, where $\hat{H}(N_{j}\mid C)=-\sum_{k=1}^{N}\log h_{\psi}^{c}(\bm{c}^{(k)})$ and $\hat{H}(N_{j}\mid C,Z)=-\sum_{k=1}^{N}\log h_{\psi}^{cz}(\bm{c}^{(k)},\bm{z}^{(k)})$.
 Here, $h_{\psi}^{c}$ and $h_{\psi}^{cz}$ are MLP classifiers trained to predict $n_{j}$ from $\bm{c}$ and $(\bm{c},\bm{z})$, respectively.
-In Table 2, we report the average value of $\frac{\hat{I}(N_{j};Z\mid C)}{H(N_{j})}$ across all $n_{j}\in\bm{n}_{y}$, which we refer to as the Uncertainty Reduction Ratio (URR).
-From these results, we draw the following conclusions: (i) regular CBMs do not consistently remove more nuisance information than VMs—in fact, contrary to common belief, CBMs can sometimes impose a weaker bottleneck than VMs; (ii) CEMs representations are those with more nuisances, i.e., they always introduce a weaker bottleneck even than VMs; and (iii) MCBMs are the models removing more nuisances from the representation, especially when $\gamma$ increases.
+In Table 3, we report the average value of $\frac{\hat{I}(N_{j};Z\mid C)}{H(N_{j})}$ across all $n_{j}\in\bm{n}_{y}$, which we call Uncertainty Reduction Ratio (URR).
+From these results, we conclude that: (i) CBMs tend to reduce nuisance information compared to VMs, though not consistently; (ii) CEMs generally preserve the largest amount of nuisance information, often exceeding even VMs; (iii) ARCBMs and HCBMs show no systematic advantage over CBMs in terms of nuisance removal; and (iv) MCBMs provide the strongest reduction of nuisance information, particularly as $\gamma$ increases, which enforces a stricter bottleneck.
 Task-unrelated information leakage
-Table 3: Average URR for task-unrelated nuisances.
+Table 4: Average value of URR for task-unrelated nuisances.
 MPI3D Shapes3D
@@
 CEM $15.5\pm 4.2$ $40.9\pm 1.8$
-MCBM (low $\gamma$) $\bm{0.0\pm 0.0}$ $\bm{0.0\pm 0.0}$
-MCBM (medium $\gamma$) $\bm{0.0\pm 0.0}$ $\bm{0.0\pm 0.0}$
-MCBM (high $\gamma$) $\bm{0.0\pm 0.0}$ $\bm{0.0\pm 0.0}$
-On the one hand, neural networks typically discard information from the input, as their layers are generally non-invertible [47, 48].
-On the other hand, there is no incentive to retain $\bm{n}_{\bar{y}}$ in the representation $\bm{z}$, since it is irrelevant for predicting $\bm{y}$.
-From this, one might expect that $\bm{z}$ should not contain information about $\bm{n}_{\bar{y}}$.
-However, prior work has shown that neural representations often retain information that is not directly relevant to the target task [24, 49].
-To investigate this phenomenon, we analyze if $\bm{n}_{\bar{y}}$ is present in $\bm{z}$ for MPI3D and Shapes3D—the only settings in our experimental setup where $\bm{n}_{\bar{y}}$ is non-empty.
-We report the URR values for the nuisance variables in $\bm{n}_{\bar{y}}$ in Table 3, and observe the following: (i) in VMs, CBMs, and CEMs, nuisance variables remain present in the representation, even when they are irrelevant to the task $\bm{y}$; and (ii) MCBMs consistently eliminate task-irrelevant nuisances across all values of $\gamma$, which is expected: since there is no incentive to preserve $\bm{n}_{\hat{y}}$, and MCBMs explicitly encourage its removal, such information tends to be discarded.
+ARCBM $8.7\pm 1.4$ $26.6\pm 0.5$
+SCBM $7.0\pm 0.3$ $21.7\pm 1.7$
+MCBM (l $\gamma$) $\bm{0.0\pm 0.0}$ $\bm{0.0\pm 0.0}$
+MCBM (m $\gamma$) $\bm{0.0\pm 0.0}$ $\bm{0.0\pm 0.0}$
+MCBM (h $\gamma$) $\bm{0.0\pm 0.0}$ $\bm{0.0\pm 0.0}$
+Neural networks typically discard input information, as their layers are non-invertible [35, 36].
+Moreover, since $\bm{n}_{\bar{y}}$ is irrelevant for predicting $\bm{y}$, there is no incentive to retain it in the representation $\bm{z}$.
+One might therefore expect $\bm{z}$ to be free of $\bm{n}_{\bar{y}}$.
+However, prior work has shown that neural representations often preserve information not directly related to the task [24, 37].
+To examine this, we analyze whether $\bm{n}_{\bar{y}}$ is present in $\bm{z}$ for MPI3D and Shapes3D—the only settings where $\bm{n}_{\bar{y}}$ is non-empty.
+Table 4 reports URR values for the nuisance variables in $\bm{n}_{\bar{y}}$ in.
+We find that: (i) as in Table 3, CEMs retain more nuisance information than even VMs; and (ii) MCBMs are the only models that consistently eliminate nuisances across all values of $\gamma$, as expected: with no incentive to preserve $\bm{n}_{\bar{y}}$ and an explicit penalty for doing so, such information is naturally discarded.
 Are concepts removed to a greater extent in MCBMs?
 Table 5: Average concepts accuracy
-CelebA CUB
-CBM $\bm{92.2\pm 0.7}$ $96.5\pm 0.1$
-CEM $92.1\pm 0.8$ $96.5\pm 0.1$
-MCBM (low $\gamma$) $92.0\pm 0.8$ $\bm{96.7\pm 0.1}$
-MCBM (medium $\gamma$) $92.1\pm 1.0$ $96.4\pm 0.1$
-MCBM (high $\gamma$) $91.8\pm 1.1$ $95.8\pm 0.4$
-One might suspect that removing nuisance information from the representation could inadvertently eliminate information about the concepts as well.
-To investigate this, we report concept prediction accuracy for CelebA and CUB in Table 4 (similar trends were observed on the remaining datasets).
-We observe no significant overall differences across models: while MCBMs show slightly lower concept accuracy on CelebA, they outperform the baselines on CUB for lower $\gamma$.
-A regular trend, however, is that increasing the hyperparameter $\gamma$ in MCBMs leads to a gradual decrease in concept accuracy, which is consistent with our expectations—stronger regularization may inadvertently suppress features that correlate with the concepts.
-These results suggest that MCBMs are effective at removing nuisance information while largely preserving the concept-relevant content in the representation.
+CIFAR-10 CUB
+CBM $84.8\pm 0.2$ $96.3\pm 0.1$
+CEM $84.8\pm 0.2$ $96.3\pm 0.1$
+ARCBM $84.3\pm 0.2$ $96.2\pm 0.1$
+SCBM $84.3\pm 0.2$ $\bm{96.5\pm 0.1}$
+MCBM (l $\gamma$) $\bm{84.9\pm 0.1}$ $96.3\pm 0.2$
+MCBM (m $\gamma$) $\bm{84.9\pm 0.2}$ $96.1\pm 0.1$
+MCBM (h $\gamma$) $84.8\pm 0.1$ $95.8\pm 0.3$
+One might worry that removing nuisance information could also inadvertently eliminate information about the concepts.
+To test this, we report concept prediction accuracy for CIFAR-10 and CUB in Table 5 (note that all models reach 100% accuracy on MPI3D and Shapes3D).
+We can observe that (i) no model consistently outperforms the others—some preserve more concept information in CIFAR-10, while others perform slightly better in CUB; and (ii) increasing $\gamma$ in MCBMs gradually reduces concept accuracy, as stronger regularization may suppress features correlated with the concepts.
+These results indicate that MCBMs effectively remove nuisance information while largely preserving concept-relevant content.
+How does this affect task performance?
+Table 6: Task accuracy
+MPI3D Shapes3D CIFAR-10 CUB
+Vanilla $\pm 0.0 100.0±0.010$$\pm 0.0$ $2.1\pm 0.6 77$$7.4\pm 0.3$
+CBM $99.9\pm 0.0$ $100.0\pm 0.0$ $3.8\pm 0.1 77$$7.6\pm 0.2$
+CEM $\pm 0.0 100.0±0.010$$\pm 0.0$ $3.1\pm 0.3 77$$7.5\pm 0.3$
+ARCBM $24.2\pm 0.6$ $29.7\pm 0.4$ $68.9\pm 0.3$ $75.5\pm 0.8$
+SCBM $24.2\pm 0.5$ $29.7\pm 0.3$ $62.3\pm 0.0$ $73.5\pm 0.4$
+MCBM (l $\gamma$) $92.7\pm 1.1$ $100.0\pm 0.0$ $72.4\pm 0.5$ $78.3\pm 0.6$
+MCBM (m $\gamma$) $46.0\pm 0.6$ $32.2\pm 1.5$ $0.8\pm 0.2 77$$7.4\pm 0.4$
+MCBM (h $\gamma$) $24.9\pm 0.3$ $30.1\pm 0.2$ $70.5\pm 0.6$ $73.5\pm 1.5$
+As previously discussed, when $\bm{n}_{y}$ is non-empty, restricting $\bm{z}$ to contain only information about $\bm{c}$ should reduce performance on $\bm{y}$: if $\bm{c}$ alone is insufficient to solve $\bm{y}$, then $\bm{z}$ cannot achieve perfect accuracy.
+In Table 6, we observe the following: (i) CBMs and CEMs reach task accuracy comparable to (or even higher than) VMs, indicating that they also rely on $\bm{n}_{y}$ to predict $\bm{y}$; (ii) ARCBMs and SCBMs achieve lower task accuracy, as they impose a bottleneck after the representations (see Appendix D); and (iii) MCBMs show decreasing task accuracy as $\gamma$ increases, reflecting the stricter bottleneck applied to the representations.
+Importantly, this reduction in task accuracy should not be viewed negatively: since all models achieve similar concept accuracy (see Table 5), it indicates that predictions rely less on nuisance information.
 4.2 Do MCBMs Yield More Interpretable Representations?
@@
 While this generally holds, we show that MCBMs yield representations that are even more interpretable.
-To support this claim, we employ two metrics: (i) Centered Kernel Alignment(CKA) [50, 51, 52] , computed between the learned representations and the concept labels (encoded as one-hot vectors), which quantifies the similarity between representation spaces using kernel methods; and (ii) Disentanglement[36], which assesses whether each dimension of the representation $z_{j}$ captures at most one concept $c_{j}$.
-We argue that these metrics offer meaningful evidence of interpretability because (i) alignment between the representations and the concepts indicates whether the information is organized in a concept-aware fashion; and (ii) disentanglement reflects the extent to which individual concepts are independently encoded in the learned representations.
-In Figure 5, we observe that: (i) although the representations learned by CBMs and CEMs exhibit substantially higher alignment with the concepts than VMs, their disentanglement levels are not reliably superior; and (ii) MCBMs consistently achieve improvements in both alignment and disentanglement.
-Thus, we conclude that removing nuisances from the representations leads to more interpretable representations.
-5.3 What Are the Pitfalls of Binarizing Multiclass Concepts?
-(a) One-vs-Rest
-(b) Multiclass
-Figure 6: Class Boundaries
-(c) One-vs-Rest
-(d) Multiclass
-Figure 7: Highest likelihood
-(a) One-vs-Rest
-(b) Multiclass
-Figure 8: Second highest likelihood
-As discussed in Section 4.1, CBMs require binarizing multiclass concepts to support interventions—i.e., they adopt a One-vs-Rest classification scheme.
-In contrast, as discussed in Section 4.4, MCBMs provide a natural and principled framework for modeling and intervening on multiclass concepts directly.
-Binarizing multiclass variables can introduce several issues, particularly in the presence of class imbalance, which is common in real-world applications.
-To illustrate this limitation, we design an experiment using a four-class spiral dataset with imbalanced class distributions.
-We train two CBMs for comparison: (i) one that binarizes concepts using a One-vs-Rest approach, and (ii) one that models concepts directly as multiclass variables.
-First, as shown in Figure 8, the One-vs-Rest model tends to favor the most frequent class more often than the Multiclass model, particularly when the true class has low prevalence and is similar (i.e., spatially close in the spiral setting) to the dominant class.
-Second, Figure 8 illustrates that, in the Multiclass model, the predicted class likelihood increases gradually from regions far from the decision boundaries toward the boundaries, resulting in a well-calibrated probability distribution.
-Such patterns are absent in the One-vs-Rest model.
-Finally, in Figure 8, we observe that the second-highest predicted likelihood in the One-vs-Rest model is close to one in certain regions—indicating that multiple classes are simultaneously assigned high confidence.
-In contrast, the Multiclass model assigns low second-highest likelihoods in regions far from the boundaries, with values increasing only near the decision boundaries.
+To support this claim, we employ two metrics: (i) Centered Kernel Alignment(CKA) [38, 39, 40], which measures the similarity between learned representations and concept labels (encoded as one-hot vectors); and (ii) Disentanglement[41], which assesses whether each dimension of the representation $z_{j}$ captures at most one concept $c_{j}$.
+These metrics capture interpretability from complementary perspectives: (i) alignment indicates whether the information is organized in a concept-aware fashion; and (ii) disentanglement reflects the extent to which individual concepts are independently encoded in the representations.
+As shown in Figure 4: (i) CBMs, CEMs, ARCBMs, and HCBMs achieve stronger alignment with concepts than Vanilla Models, yet they do not consistently improve disentanglement; and (ii) MCBMs consistently improve both alignment and disentanglement, particularly as $\gamma$ increases, reflecting the removal of additional nuisances.
+These results suggest that explicitly eliminating nuisances leads to more interpretable representations.
+4.3 Do MCBMs Enable More Reliable Interventions?
+(a) CIFAR-10
+(b) CUB
+Figure 5: Error (y-axis) versus percentage of concepts intervened (x-axis) across different models.
+Another key property often attributed to CBMs is their capacity to support interventions.
+Although standard CBM intervention methods suffer from theoretical limitations (see Section 5), they can still be applied in practice.
+Accordingly, for CBMs we follow the procedure in Section 2.3, for ARCBMs and SCBMs we use the procedure described in Appendix D, and for MCBMs we apply the method in Section 2.4.
+To evaluate intervention reliability across models, we adopt the standardized protocol of Koh et al.
+[9], which tracks how prediction error changes as the number of intervened concepts increases.
+Following Shin et al.
+[42], we intervene on the concepts predicted with the lowest confidence.
+Results for CIFAR-10 and CUB are shown in Figure 5 (note that interventions have minimal effect on MPI3D and Shapes3D, since—as shown in Table 6—the concepts are not strong predictors of the task).
+From these results, four main observations emerge: (i) CBMs may even increase error when multiple concepts are intervened—an effect of nuisance information leaking into the representation due to the absence of a proper bottleneck; (ii) ARCBMs and SCBMs mitigate this by applying an Information Bottleneck after the representations, which improves intervention effectiveness while leaving interpretability unchanged (Figure 4); (iii) in MCBMs, intervention gains remain largely invariant to $\gamma$, as indicated by the nearly parallel trends across its different values; and (iv) at low and medium $\gamma$, MCBMs deliver the strongest intervention performance, especially when fewer than 100% of the concepts are intervened, clearly outperforming ARCBMs and SCBMs.
+5 Other Fundamental Theoretical Flaws of CBMs
+As briefly discussed in Section 2.3, CBMs—unlike MCBMs—do not provide a principled mechanism to estimate $p(z_{j}\mid c_{j})$.
+To enable interventions, two assumptions are typically introduced: (i) multiclass concepts are handled using a One-vs-Rest scheme, and (ii) interventions are implemented via the sigmoid inverse function, i.e., $p(z_{j}\mid c_{j})\approx\delta\left(z_{j}-\sigma^{-1}(c_{j})\right)$.
+These assumptions are not only ad hoc but also theoretically incorrect, as we explain and illustrate with toy experiments below.
+One-vs-Rest Limitations
+Figure 6: CBM (top) versus MCBM (bottom): class boundaries (left), highest predicted likelihood (middle), and second-highest predicted likelihood (right).
+One-vs-Rest strategies exhibit several limitations: (i) individual binary classifiers tend to be biased toward the negative class, and (ii) the predicted probabilities across classifiers are typically uncalibrated [43].
+To illustrate these issues, we design an experiment where the concepts are defined based on a four-class spiral dataset with imbalanced class distributions.
+We train: (i) a CBM with One-vs-Rest binarization, and (ii) an MCBM modeling concepts directly as multiclass variables.
+The results reveal three major shortcomings of CBMs trained with the One-vs-Rest strategy.
+First, Figure 6 (left) shows that they overpredict the most frequent class, especially when the true class is rare and spatially close to the dominant.
+Second, Figure 6 (middle) shows that CBMs produce poorly calibrated predictions, lacking the smooth likelihood transitions observed in MCBMs.
+Finally, Figure 6 (right) shows that CBMs often assign near-one likelihoods to multiple classes simultaneously, whereas MCBMs confine high secondary likelihoods to regions of overlap, producing more reliable uncertainty estimates.
+Importantly, these effects may not be captured by standard metrics like accuracy, yet they represent fundamental flaws from a probabilistic perspective.
+Sigmoid Inverse Function
+Figure 7: Density (y-axis) vs. $z$ (x-axis)
+[…]
 6 Conclusions
-In this paper, we argue that—contrary to common belief—Concept Bottleneck Models (CBMs) do not enforce a true bottleneck: while representations are encouraged to retain concept-related information, they are not constrained to discard nuisance information.
-This limitation results in a lack of theoretical guarantees for intervention procedures and reduced interpretability in the representation space.
+In this paper, we argue that—contrary to common belief—Concept Bottleneck Models (CBMs) do not enforce a true bottleneck: although representations are encouraged to retain concept-related information, they are not constrained to discard nuisance information.
+This limitation undermines interpretability and provides no theoretical guarantees for intervention procedures.
 To address this, we propose Minimal Concept Bottleneck Models (MCBMs), which introduce an Information Bottleneck in the representation space via an additional loss term derived through variational approximations.
-Beyond establishing a proper bottleneck, our formulation ensures that interventions do not violate Bayesian principles and introduces greater design flexibility compared to standard CBMs.
-Through empirical analysis, we show that CBMs—and their existing variants—fail to effectively remove nuisance information from the representation, even when such information is irrelevant to the target task.
-In contrast, MCBMs successfully eliminate nuisances while preserving concept-relevant information, resulting in more interpretable representations and principled intervention mechanisms.
-Finally, we highlight a common limitation of CBMs: the need to binarize multiclass concepts due to architectural constraints.
-MCBMs natively support multiclass concepts, thereby overcoming the drawbacks introduced by this binarization process
+Beyond enforcing a proper bottleneck, our formulation ensures that interventions remain consistent with Bayesian principles.
+Empirically, we show that CBMs and their variants fail to remove non-concept information from the representation, even when irrelevant to the target task.
+In contrast, MCBMs effectively eliminate such information while preserving concept-relevant content, yielding more interpretable representations and principled interventions.
+Finally, we highlight fundamental flaws in the intervention process of CBMs, which MCBMs overcome due to their principled design.
 Acknowledgments and Disclosure of Funding
@@
 Inputs $\bm{x}$ are defined by some concepts $\{c_{j}\}_{j=1}^{m}$ and nuisances $\bm{n}$; and targets $\bm{y}$ are defined by $\bm{x}$ (gray arrows).
-Vanilla models obtain the representations $\{z_{j}\}_{j=1}^{m}$ from $\bm{x}$ through the encoder $p_{\theta}(z|x)$ (black arrows) and solve the task $\hat{y}$ sequentially through the task head $q_{\phi}(y|z)$ (green arrows).
-Concept Bottleneck Models try to make a prediction $\hat{c}_{j}$ of each concept $c_{j}$ from one representation $z_{j}$ through the concept head $q(\hat{c}_{j}|z_{j})$ (blue arrows).
-Minimal CBMs (ours) try to make a prediction $\hat{z}_{j}$ of each representation $z_{j}$ from one concept $c_{j}$ through the representation head $q(\hat{z}_{j}|c_{j})$ (red arrows).
+Vanilla models obtain the representations $\{z_{j}\}_{j=1}^{m}$ from $\bm{x}$ through the encoder $\color[rgb]{0,1,1}p_{\theta}(\bm{z}|\bm{x})$ (cyan arrows) and solve the task $\hat{\bm{y}}$ sequentially through the task head $\color[rgb]{0,1,0}q_{\phi}(\hat{\bm{y}}|\bm{z})$ (green arrows).
+Concept Bottleneck Models make a prediction $\hat{c}_{j}$ of each concept $c_{j}$ from one representation $z_{j}$ through the concept head $\color[rgb]{0,0.49609375,1}q(\hat{c}_{j}|z_{j})$ (blue arrows).
+Minimal CBMs make a prediction $z_{j}$ of each representation $z_{j}$ from one concept $c_{j}$ through the representation head $\color[rgb]{1,0,0}q(z_{j}|c_{j})$ (red arrows).
 Appendix B Details of Derivations
@@
 16:   Encode: Compute $\mu_{\theta}^{(k)}\leftarrow f_{\theta}\left(x^{(k)}\right)$
-17:         Sample noise $\epsilon\sim\mathcal{N}(0,I)$triangleright Reparameterization trick with only one sample
+17:   Sample noise $\epsilon\sim\mathcal{N}(0,I)$ $\triangleright$ Reparameterization trick with only one sample
 18:   Reparameterize: $\bm{z}^{(k)}\leftarrow\mu_{\theta}^{(k)}+\sigma_{x}\odot\epsilon$
-19:         Task prediction: $\hat{\bm{y}}^{(k)}\leftarrow g^{y}_{\phi}\left(\bm{z}^{(k)}\right)$triangleright Similar to VMs
+19:   Task prediction: $\hat{\bm{y}}^{(k)}\leftarrow g^{y}_{\phi}\left(\bm{z}^{(k)}\right)$ $\triangleright$ Similar to VMs
 20:   Task loss: $\mathcal{L}_{y}^{(k)}\leftarrow\left\|\bm{y}^{(k)}-\hat{\bm{y}}^{(k)}\right\|^{2}\text{ if }\bm{y}\text{ is continuous else }\text{CE}\left(\bm{y}^{(k)},\bm{\hat{y}}^{(k)}\right)$
 21:   for all $j=1,\dots,n$ do
-22:              Concept $j$ prediction: $\hat{c}^{(k)}_{j}\leftarrow g^{c}_{\phi,j}\left(z^{(k)}_{j}\right)$triangleright Similar to CBMs
+22:     Concept $j$ prediction: $\hat{c}^{(k)}_{j}\leftarrow g^{c}_{\phi,j}\left(z^{(k)}_{j}\right) \triangleright$ Similar to CBMs
 23:     Concept $j$ loss: $\mathcal{L}_{c,j}^{(k)}\leftarrow\left\|c^{(k)}_{j}-\hat{c}^{(k)}_{j}\right\|^{2}\text{ if }c_{j}\text{ is continuous else }\text{CE}\left(c^{(k)}_{j},\hat{c}^{(k)}_{j}\right)$
-24:              Representation $j$ prediction: $\hat{z}^{(k)}_{j}\leftarrow g^{z}_{j}\left(c^{(k)}_{j}\right)$triangleright Novelty in MCBMs
+24:     Representation $j$ prediction: $\hat{z}^{(k)}_{j}\leftarrow g^{z}_{j}\left(c^{(k)}_{j}\right) \triangleright$ Novelty in MCBMs
 25:     Representation $j$ loss: $\mathcal{L}_{z,j}^{(k)}\leftarrow\left\|z^{(k)}_{j}-\hat{z}^{(k)}_{j}\right\|^{2}$
@@
 30:end while
-Appendix D Experiments Details
-D.1 Hyperparameters for Section 5
+Appendix D Hard Concept Bottleneck Models
+As discussed in Section 3, one of the most successful approaches to mitigating information leakage is the family of Hard Concept Bottleneck Models (HCBMs) [18].
+Unlike MCBMs, which enforce an Information Bottleneck (IB) directly at the representation level $\bm{z}$, HCBMs operate by pruning the information contained in $\bm{z}$ before producing the task prediction $\hat{\bm{y}}$.
+In doing so, they encourage $\hat{\bm{y}}$ to depend solely on $\bm{c}$, thereby yielding interventions that are more reliable in practice.
+This approach differs fundamentally from MCBMs: whereas Equation 4 in MCBMs minimizes over $Z_{j}$, thereby removing nuisances directly from the representation $z_{j}$, HCBMs optimize at the prediction level, focusing on $\hat{Y}$.
+Formally, they are trained to minimize:
+$\vskip 2.0pt\min_{\hat{Y}}I(\hat{Y};X|C)=\min_{\theta,\phi}\mathbb{E}_{p(x,c_{j})}\left[D_{KL}\left(p_{\theta,\phi}(\hat{\bm{y}}|x)||p(\hat{\bm{y}}|c)\right)\right]$ (24)
+That is, the optimization is performed directly over the predictions $\hat{\bm{y}}$. To implement this, it defines:
+$p_{\theta,\phi}(\hat{\bm{y}}|x)=\iiint q_{\phi}\left(\hat{\bm{y}}|\hat{\bm{c}}^{b}\right){\color[rgb]{1,0,1}p\left(\hat{\bm{c}}^{b}|\hat{\bm{c}}\right)}q\left(\hat{\bm{c}}|\bm{z}\right)p_{\theta}\left(\bm{z}|\bm{x}\right)\,d\hat{\bm{c}}^{b}\,d\hat{\bm{c}}\,d\bm{z}$ (25)
+where $q_{\phi}\left(\hat{\bm{y}}\mid\hat{\bm{c}}^{b}\right)$ denotes the new task head, $q\left(\hat{\bm{c}}\mid\bm{z}\right)$ the concept head, and $p_{\theta}\left(\bm{z}\mid\bm{x}\right)$ the encoder, as defined in Section 2.
+The distribution
+$p\left(\hat{c}_{j}^{b}|\hat{c}_{j}\right)=\delta\left(\hat{c}_{j}^{b}-\Theta\left(\hat{c}_{j}-0.5\right)\right)$ (26)
+referred to as the binarizing head, applies the Heaviside step function $\Theta$ to produce a binary version $\hat{c}^{b}_{j}\in\{0,1\}$ of $\hat{c}_{j}$.
+In this way, HCBMs enforce an ad-hoc Information Bottleneck by predicting $\hat{\bm{y}}$ from binarized concept representations.
+This process is schematized in Figure 9.
+How are Interventions Performed in HCBMs?
+Because HCBMs introduce a bottleneck immediately before predicting $\hat{\bm{y}}$, the intervention process is more straightforward than in standard CBMs.
+More specifically, interventions are carried out according to:
+$p(\hat{\bm{y}}|c_{j}=\alpha,\bm{x})=\iint q_{\phi}(\hat{\bm{y}}|\hat{c}^{b}_{j},\bm{\hat{c}}^{b}_{\backslash j})p(\hat{c}^{b}_{j}|c_{j}=\alpha)p_{\theta}(\bm{\hat{c}}^{b}_{\backslash j}|\bm{x})\,d\hat{c}^{b}_{j}\,d\bm{\hat{c}}^{b}_{\backslash j}$ (27)
+Here, $p(\hat{c}^{b}_{j}|c_{j}=\alpha)$ is approximated as $\delta(\hat{c}^{b}_{j}-\alpha)$, while the other distributions are available in closed form.
+Although this procedure provides stronger guarantees than the intervention mechanisms in standard CBMs, two main issues remain:
+  1. (i)
+    The optimization is performed over the predictions $\hat{\bm{y}}$ instead of the representation $\bm{z}$.
+As a result, the representations themselves are not necessarily interpretable, as evidenced in Figure 4, which limits one of the core motivations for adopting CBMs in the first place.
+  2. (ii)
+    HCBMs still require binarizing multiclass concepts, which introduces theoretical limitations and practical drawbacks, as further discussed in Section 5.
+Figure 9: Graphical models of HCBMs with two concepts and two-dimensional representations.
+Hard Concept Bottleneck Models obtain a binarized version $\hat{c}^{b}_{j}$ of each predicted concept $\hat{c}_{j}$ through the binarizing head $\color[rgb]{1,0,1}p(\hat{c}^{b}_{j}\mid\hat{c}j)$ (fuchsia arrows).
+Unlike the models in Figure 2, HCBMs predict the task output $\hat{\bm{y}}$ from the binarized concepts $\hat{c}^{b}_{j}$ using the new task head $\color[rgb]{0,1,0}q_{\phi}(\hat{\bm{y}}\mid\bm{z})$ (green arrows).
+Appendix E Experiments Details
+E.1 Hyperparameters for Section 4
 Table 7: Hyperparameters for Section 4.
@@
 We emphasize that CBMs are, by design, restricted to use invertible $g^{c}_{\phi}$ to enable intervention procedures.
-CIFAR-10 MPI3D Shapes3D CelebA CUB
-$f_{\theta}$ architecture ResNet20 ResNet20 ResNet20 ResNet18 InceptionV3
-$f_{\theta}$ pretraining None None None ImageNet ImageNet
-$g^{y}_{\phi}$ hidden layers 64 64 64 64 256
-low $\gamma$ 1 1 1 0.05 0.1
-medium $\gamma$ 5 3 3 0.1 0.3
-high $\gamma$ 10 5 5 0.3 0.5
-number of epochs 50 50 50 120 250
-batch size 128 128 128 128 128
-optimizer SGD SGD SGD SGD SGD
-learning rate $6\times 10^{-3}$ $6\times 10^{-3}$ $6\times 10^{-3}$ $2\times 10^{-2}$ $2\times 10^{-2}$
-momentum 0.9 0.9 0.9 0.9 0.9
-weight decay $4\times 10^{-5}$ $4\times 10^{-5}$ $4\times 10^{-5}$ $4\times 10^{-5}$ $4\times 10^{-5}$
-scheduler Step Step Step Step Step
-step size (epochs) 20 20 20 50 100
-scheduler $\gamma$ 0.1 0.1 0.1 0.1 0.1
-D.2 Datasets
-CIFAR-10
-The CIFAR-10 dataset consists of 60,000 color images of size 32×32, evenly distributed across 10 object classes, with 50,000 training and 10,000 test samples.
-Among these, 6 classes correspond to animals and 4 to vehicles.
-We define the target task $\bm{y}$ as a binary classification between animals and vehicles.
-The subset $\bm{n}_{y}$ includes one vehicle and one animal class—specifically, airplane and bird—while $\bm{c}$ comprises the remaining 8 object classes.
+MPI3D Shapes3D CIFAR-10 CUB
+$f_{\theta}$ architecture ResNet20 ResNet20 2 conv. layers InceptionV3
+$f_{\theta}$ pretraining None None None ImageNet
+$g^{y}_{\phi}$ hidden layers 64 64 64 256
+low $\gamma$ 1 1 0.1 0.05
+medium $\gamma$ 3 3 0.3 0.1
+high $\gamma$ 5 5 0.5 0.3
+number of epochs 50 50 200 250
+batch size 128 128 128 128
+optimizer SGD SGD Adam SGD
+learning rate $6\times 10^{-3}$ $6\times 10^{-3}$ $1\times 10^{-4}$ $2\times 10^{-2}$
+momentum 0.9 0.9 0. 0.9
+weight decay $4\times 10^{-5}$ $4\times 10^{-5}$ $4\times 10^{-5}$ $4\times 10^{-5}$
+scheduler Step Step Step Step
+step size (epochs) 20 20 80 100
+scheduler $\gamma$ 0.1 0.1 0.1 0.1
+E.2 Datasets
 MPI3D
 This is a synthetic dataset with controlled variation across seven generative factors: object shape, object color, object size, camera height, background color, horizontal axis, and vertical axis.
-In our setup, $\bm{y}$ corresponds to the object shape, $\bm{n}_{y}$ to the horizontal axis, $\bm{n}{\bar{y}}$ to the vertical axis, and $\bm{c}$ to the remaining generative factors.
-To ensure consistency in the mapping between concepts and the target, we filter the dataset such that any combination of elements in $\left\{\bm{c},\bm{n}_{y}\right\}$ corresponds to a unique value of $\bm{y}$.
+In our setup, $\bm{y}$ corresponds to the object shape, $\bm{n}_{y}$ to the horizontal axis, $\bm{n}_{\bar{y}}$ to the vertical axis, and $\bm{c}$ to the remaining generative factors.
+To ensure consistency in the mapping between concepts and task nuisances and the target, we filter the dataset such that any combination of elements in $\left\{\bm{c},\bm{n}_{y}\right\}$ corresponds to a unique value of $\bm{y}$.
 All invalid combinations are removed accordingly.
@@
 We follow the same filtering strategy as in MPI3D to construct this configuration: we retain only those samples for which each combination of $\left\{\bm{c},\bm{n}_{y}\right\}$ uniquely determines $\bm{y}$, removing all invalid configurations.
-CelebA
-: The CelebA dataset consists of over 200,000 celebrity face images, each annotated with 40 binary facial attributes.
-In our setup, the task variable $\bm{y}$ corresponds to binary gender classification.
-For each run, eight attributes are randomly selected as the concept set $\bm{c}$, while the remaining attributes are treated as nuisances $\bm{n}_{y}$.
-Since most attributes exhibit some degree of correlation with gender, we set $\bm{n}_{\bar{y}}$ to the empty set.
+CIFAR-10
+CIFAR-10 is a widely used image classification benchmark consisting of 60,000 natural images of size 32 $\times$ 32, divided into 10 classes (e.g., airplanes, automobiles, birds, cats, etc.).
+The dataset is split into 50,000 training and 10,000 test images, with balanced class distributions.
+To reduce the need for manual concept annotations, the concepts are synthetically derived following the methodology of [23].
+A total of 143 attributes are extracted using GPT-3 [33]; 64 form the concept set $\bm{c}$, while the rest define the nuisance set $\bm{n}_{y}$.
+Binary values are obtained with the CLIP model [44] by comparing the similarity of each image to the embedding of an attribute and to its negative counterpart.
 CUB
-: The Caltech-UCSD Birds (CUB) dataset contains 11,788 images of 200 bird species, annotated with part locations, bounding boxes, and 312 binary attributes.
-Following the approach of [9], we retain only the attributes that are present in at least 10 species (based on majority voting), resulting in a filtered set of 112 attributes.
+The Caltech-UCSD Birds (CUB) dataset contains 11,788 images of 200 bird species, annotated with part locations, bounding boxes, and 312 binary attributes.
+Following the approach of Koh et al.
+[9], we retain only the attributes that are present in at least 10 species (based on majority voting), resulting in a filtered set of 112 attributes.
 These attributes are grouped into 27 semantic clusters, where each group is defined by a common prefix in the attribute names.
 In our setup, the task variable $\bm{y}$ is to classify the bird species.
-The concept set $\bm{c}$ consists of the attributes belonging to 7 randomly selected groups (per run), while the nuisance set $\bm{n}y$ includes the attributes from the remaining 20 groups.
-Since most attributes exhibit some correlation with the classification task, we set $\bm{n}{\bar{y}}$ to the empty set.
+The concept set $\bm{c}$ consists of the attributes belonging to 12 randomly selected groups (per run), while the nuisance set $\bm{n}y$ includes the attributes from the remaining 20 groups.
+Since most attributes exhibit some correlation with the classification task, we set $\bm{n}_{\bar{y}}$ to the empty set.

### v3

@@
 1 Introduction
-Most machine learning models operate by learning data representations—compressed versions of the input that retain the essential information needed to solve a given task [1].
+Most machine learning models operate by learning data representations—compressed versions of the input that retain the essential information needed to solve a given task (Bengio et al., 2013).
 However, these representations often encode information in ways that are not easily interpretable by humans.
-This lack of interpretability becomes especially problematic in sensitive domains such as healthcare [2, 3, 4], finance [5, 6], and autonomous driving [7, 8].
-To address this issue, Concept Bottleneck Models (CBMs) have been proposed, which enforce representations to be defined in terms of a set of human-understandable concepts [9].
-Formally, given a task $\bm{y}$ and a set of concepts $\bm{c}=\{c_{j}\}_{j=1}^{m}$, Vanilla Models (VMs) are trained with a single objective: their representations $\bm{z}$ should capture the information necessary to predict $\bm{y}$ accurately.
-CBMs extend this framework by adding a second objective: each concept $c_{j}$ must be recoverable from a designated component $z_{j}\in\bm{z}$.
-By enforcing this additional constraint, CBMs are claimed to offer: (i) improved interpretability of the representation space, and (ii) the ability to intervene on specific concepts by modifying $z_{j}$ and propagating these changes to the predictions.
-However, CBMs are prone to a phenomenon known as information leakage [10, 11], where the representation $\bm{z}$ encodes input information that cannot be attributed to the predefined concepts $\bm{c}$.
+This lack of interpretability becomes especially problematic in sensitive domains such as healthcare (Ahmad et al., 2018; Xie et al., 2020; Jin et al., 2022), finance (Brigo et al., 2021; Liu et al., 2023), and autonomous driving (Kim and Canny, 2017; Xu et al., 2024b).
+To address this issue, Concept Bottleneck Models (CBMs) have been proposed, which enforce representations to be defined in terms of a set of human-understandable concepts (Koh et al., 2020).
+Given an input $\bm{x}$ and target $\bm{y}$, Vanilla Models (VMs) are trained with a single goal: the representation $\bm{z}$ derived from $\bm{x}$ should encode all information needed to accurately predict $\bm{y}$.
+In many settings, additional side information—often referred to as concepts—is available, denoted by $\bm{c}=\{c_{j}\}_{j=1}^{m}$.
+Concept Bottleneck Models (CBMs) leverage this by extending VMs with a second objective: each concept $c_{j}$ must be recoverable from a designated component $z_{j}\in\bm{z}$.
+By enforcing this additional constraint, CBMs are purported to provide: (i) enhanced interpretability of the learned representation space, and (ii) the ability to perform targeted interventions on specific concepts by manipulating $z_{j}$ and propagating the resulting changes to the model’s predictions.
+However, CBMs are prone to a phenomenon known as information leakage (Margeloiu et al., 2021; Mahinpei et al., 2021), where the representation $\bm{z}$ encodes input information that cannot be attributed to the predefined concepts $\bm{c}$.
 We refer to this additional information as nuisances $\bm{n}$.
 Information leakage raises two main concerns: (i) it undermines interpretability, since $z_{j}$ cannot be fully explained by its corresponding concept $c_{j}$; and (ii) it compromises the validity of interventions—modifying $z_{j}$ may alter not only the associated concept $c_{j}$, but also other unintended information encoded in $z_{j}$.
-We argue that information leakage stems from a fundamental limitation in the current formulation of CBMs: the absence of an explicit Information Bottleneck (IB) [12] that actively constrains $z_{j}$ to exclude information unrelated to $c_{j}$.
+We argue that information leakage stems from a fundamental limitation in the current formulation of CBMs: the absence of an explicit Information Bottleneck (IB) (Tishby et al., 2000) that actively constrains $z_{j}$ to exclude information unrelated to $c_{j}$.
 While the second objective in CBMs encourages each $z_{j}$ to retain $c_{j}$ in its entirety, it does not enforce that $z_{j}$ captures only information about $c_{j}$.
@@
 This ensures that $z_{j}$ not only retains all the information about its associated concept $c_{j}$, but also excludes any information unrelated to $c_{j}$, as summarized in Table 1.
-The name reflects that $z_{j}$ is trained to be a minimal sufficient statistic of $c_{j}$ [13, 14], in contrast to traditional CBMs where $z_{j}$ is optimized to be merely a sufficient statistic of $c_{j}$.
+The name reflects that $z_{j}$ is trained to be a minimal sufficient statistic of $c_{j}$ (Fisher, 1922; 1935), in contrast to traditional CBMs where $z_{j}$ is optimized to be merely a sufficient statistic of $c_{j}$.
 As illustrated in Figure 1, this design yields disentangled representations that directly address the two shortcomings of CBMs: (i) it improves interpretability, since $z_{j}$ can be fully explained by its corresponding concept $c_{j}$; and (ii) it enables valid interventions—modifying $z_{j}$ affects only the associated concept $c_{j}$.
@@
 2.1 Data Generative Process
-For this scenario, we consider inputs $\bm{x}\in\mathcal{X}$, targets $y\in\mathcal{Y}$, concepts $\bm{c}=\{c_{j}\}_{j=1}^{m}\in\mathcal{C}$ and nuisances $\bm{n}\in\mathcal{N}$, such that $p(\bm{x},\bm{y},\bm{c},\bm{n})=p(\bm{x}|\bm{c},\bm{n})p(\bm{y}|\bm{x})$, i.e., the inputs $\bm{x}$ are described by the concepts $\bm{c}$ and the nuisances $\bm{n}$, and the targets $\bm{y}$ are fully described by the input $\bm{x}$.
+For this scenario, we consider inputs $\bm{x}\in\mathcal{X}$, targets $\bm{y}\in\mathcal{Y}$, concepts $\bm{c}=\{c_{j}\}_{j=1}^{m}\in\mathcal{C}$ and nuisances $\bm{n}\in\mathcal{N}$, such that $p(\bm{x},\bm{y},\bm{c},\bm{n})=p(\bm{x}|\bm{c},\bm{n})p(\bm{y}|\bm{x})p(\bm{c},\bm{n})$, i.e., the inputs $\bm{x}$ are described by the concepts $\bm{c}$ and the nuisances $\bm{n}$, and the targets $\bm{y}$ are fully described by the input $\bm{x}$.
 The only difference between $\bm{c}$ and $\bm{n}$ is that the former are observed, while the latter are not or, in other words, labels on $\bm{c}$ are provided to us.
@@
 We refer to models trained to address this problem as Vanilla Models (VMs).
-These models typically operate by first extracting an intermediate representation $\bm{z}\in\mathcal{Z}$ from the input $\bm{x}$ via an encoder $p_{\theta}(\bm{z}\mid\bm{x})$.
-Subsequently, a prediction $\hat{y}\in\mathcal{Y}$ is produced from $\bm{z}$ using a task head $q_{\phi}(\hat{\bm{y}}\mid\bm{z})$.
+These models typically operate by first extracting an intermediate representation $\bm{z}\in\mathcal{Z}$ from the input $\bm{x}$ via an encoder $p_{\theta}(\bm{z}\mid\bm{x})$ parameterized by $\theta$.
+Subsequently, a prediction $\hat{y}\in\mathcal{Y}$ is produced from $\bm{z}$ using a task head $q_{\phi}(\hat{\bm{y}}\mid\bm{z})$ parameterized by $\phi$.
 Since $\bm{z}$ is intended to facilitate accurate prediction of $\bm{y}$, the mutual information between $\bm{z}$ and $\bm{y}$, denoted $I(Z;Y)$, should be maximized.
@@
 Black edges represent the encoder $p_{\theta}(\bm{z}\mid\bm{x})$, while green edges indicate the task head $q_{\phi}(\hat{\bm{y}}\mid\bm{z})$.
-The encoder is typically chosen to be deterministic, i.e., $p_{\theta}(\bm{z}\mid\bm{x})=\delta\left(\bm{z}-f_{\theta}(\bm{x})\right)$.
+The encoder is typically chosen to be deterministic, i.e., $p_{\theta}(\bm{z}\mid\bm{x})=\delta\left(\bm{z}-f_{\theta}(\bm{x})\right)$, where $f_{\theta}:\mathcal{X}\to\mathcal{Z}$ is a neural-network–parameterized mapping.
 However, for tractability reasons (see Section 2.4), we adopt a stochastic formulation where $p_{\theta}(\bm{z}\mid\bm{x})=\mathcal{N}\left(\bm{z};f_{\theta}(\bm{x}),\sigma_{x}^{2}I\right)$, as summarized in Table 2.
@@
   1. (i)
-    Concepts $c_{j}\in\bm{c}$, are assumed to be binary. If a concept is originally multiclass with $k$ categories, it is converted into $k$ binary concepts (One-vs-Rest [15]).
+    Concepts $c_{j}\in\bm{c}$, are assumed to be binary. If a concept is originally multiclass with $k$ categories, it is converted into $k$ binary concepts (One-vs-Rest (Rifkin and Klautau, 2004)).
   2. (ii)
@@
 However, this is ill-defined at the binary extremes, as $\sigma^{-1}(1)=-\sigma^{-1}(0)=\infty$.
-To address this, in practice, $p(z_{j}\mid c_{j}=0)$ and $p(z_{j}\mid c_{j}=1)$ are set as the 5th and 95th percentiles of the empirical distribution of $z_{j}$, respectively [9].
+To address this, in practice, $p(z_{j}\mid c_{j}=0)$ and $p(z_{j}\mid c_{j}=1)$ are set as the 5th and 95th percentiles of the empirical distribution of $z_{j}$, respectively (Koh et al., 2020).
 This workaround, however, introduces two crucial issues discussed in Section 5.
 Table 2: Distributions considered in this work.
-$f_{\theta}$ is typically modeled by a large neural network while $g^{y}_{\phi}$, $g^{c}_{\phi}$ and $g^{z}_{\phi}$ by simpler neural networks.
-We consider $\bm{z}$ to be always continuous.
+$f_{\theta}:\mathcal{X}\to\mathcal{Z}$ is typically a large neural network while $g^{y}_{\phi}:\mathcal{Z}\to\mathcal{Y}$, $g^{c}_{\phi}:\mathcal{Z}\to\mathcal{C}$ and $g^{z}_{\phi}:\mathcal{C}\to\mathcal{Z}$ are comparatively lightweight networks (see Appendix F.1).
+Throughout this work, we model $\bm{z}$ as a continuous latent representation.
 $p_{\theta}(z|x)$ $q_{\phi}(\hat{y}|z)$ $q_{\phi}(\hat{c}_{j}|z_{j})$ $q_{\phi}(\hat{z}_{j}|c_{j})$
@@
     For categorical concepts, $g^{z}_{\phi}(c_{j})=\lambda\cdot\mathrm{one\_hot}(c_{j})$.
-This mirrors the Prototypical Learning [16] approach where class-dependent prototypes $\{g^{z}_{\phi}(c_{j})\}_{j=1}^{m}$ are fixed.
+This mirrors the Prototypical Learning (Snell et al., 2017) approach where class-dependent prototypes $\{g^{z}_{\phi}(c_{j})\}_{j=1}^{m}$ are fixed.
   3. (iii)
@@
 First, we approximate the expectations over $p(\bm{x},\bm{y})$ and $p(\bm{x},c_{j})$ using the empirical data distribution, replacing integrals with summations over the dataset.
-Second, to enable gradient-based optimization through the stochastic encoder, we apply the reparameterization trick [17]: $\mathbb{E}_{p_{\theta}(\bm{z}|\bm{x})}\left[\log q_{\phi}(\bm{y}|\bm{z})\right]\approx\sum_{i}\log q_{\phi}\left(\bm{y}|f^{\prime}_{\theta}\left(x,\epsilon^{(i)}\right)\right)$ and $E_{p_{\theta}(z_{j}|\bm{x})}\left[\log q_{\phi}(\hat{c}_{j}|z_{j})\right]\approx\sum_{i}\log q_{\phi}\left(\hat{c}_{j}\big|f^{\prime}_{\theta,j}\left(\bm{x},\epsilon^{(i)}\right)\right)$, where $f^{\prime}_{\theta}(\bm{x},\epsilon)=f_{\theta}(\bm{x})+\sigma^{2}_{x}I\epsilon$ (due to the choice of $p_{\theta}(\bm{z}|\bm{x})$ in Table 2), $\epsilon\sim\mathcal{N}\left(0,I\right)$ and $f^{\prime}_{\theta,j}\left(x,\epsilon\right)$ corresponds to the element $j$ of $f^{\prime}_{\theta}\left(\bm{x},\epsilon\right)$.
+Second, to enable gradient-based optimization through the stochastic encoder, we apply the reparameterization trick (Kingma, 2013): $\mathbb{E}_{p_{\theta}(\bm{z}|\bm{x})}\left[\log q_{\phi}(\bm{y}|\bm{z})\right]\approx\sum_{i}\log q_{\phi}\left(\bm{y}|f^{\prime}_{\theta}\left(x,\epsilon^{(i)}\right)\right)$ and $E_{p_{\theta}(z_{j}|\bm{x})}\left[\log q_{\phi}(\hat{c}_{j}|z_{j})\right]\approx\sum_{i}\log q_{\phi}\left(\hat{c}_{j}\big|f^{\prime}_{\theta,j}\left(\bm{x},\epsilon^{(i)}\right)\right)$, where $f^{\prime}_{\theta}(\bm{x},\epsilon)=f_{\theta}(\bm{x})+\sigma^{2}_{x}I\epsilon$ (due to the choice of $p_{\theta}(\bm{z}|\bm{x})$ in Table 2), $\epsilon\sim\mathcal{N}\left(0,I\right)$ and $f^{\prime}_{\theta,j}\left(x,\epsilon\right)$ corresponds to the element $j$ of $f^{\prime}_{\theta}\left(\bm{x},\epsilon\right)$.
 Combining the considerations above yields the final objective for training MCBMs, as shown in Equation 6, where $\beta$ and $\gamma$ are hyperparameters.
@@
 Figure 2: Graphical models of the different systems described for two concepts and two-dimensional representations.
-We provide in Appendix A the analogous figure for $m$ concepts and $m$-dimensional representations.
+Appendix A shows the analogous figure for $m$ concepts and $m$-dimensional representations.
 Inputs $\bm{x}$ are defined by some concepts $\{c_{j}\}_{j=1}^{m}$ and nuisances $\bm{n}$; and targets $\bm{y}$ are defined by $\bm{x}$ (gray arrows).
 Vanilla models obtain the representations $\{z_{j}\}_{j=1}^{m}$ from $\bm{x}$ through the encoder $\color[rgb]{0,1,1}p_{\theta}(\bm{z}|\bm{x})$ (cyan arrows) and solve the task $\hat{\bm{y}}$ sequentially through the task head $\color[rgb]{0,1,0}q_{\phi}(\hat{\bm{y}}|\bm{z})$ (green arrows).
-Concept Bottleneck Models make a prediction $\hat{c}_{j}$ of each concept $c_{j}$ from one representation $z_{j}$ through the concept head $\color[rgb]{0,0.49609375,1}q(\hat{c}_{j}|z_{j})$ (blue arrows).
-Minimal CBMs make a prediction $z_{j}$ of each representation $z_{j}$ from one concept $c_{j}$ through the representation head $\color[rgb]{1,0,0}q(z_{j}|c_{j})$ (red arrows).
+Concept Bottleneck Models make a prediction $\hat{c}_{j}$ of each concept $c_{j}$ from one representation $z_{j}$ through the concept head $\color[rgb]{0,0.49609375,1}q_{\phi}(\hat{c}_{j}|z_{j})$ (blue arrows).
+Minimal CBMs make a prediction $\hat{z}_{j}$ of each representation $z_{j}$ from one concept $c_{j}$ through the representation head $\color[rgb]{1,0,0}q_{\phi}(\hat{z}_{j}|c_{j})$ (red arrows).
 3 Related Work
 Information Leakage in Concept Bottleneck Models
-Information leakage occurs when the learned representation $\bm{z}$ encodes information outside the concept set $\bm{c}$, reducing both interpretability and intervenability [10, 11].
-This arises when the Markovian assumption fails—i.e., when the target $\bm{y}$ is not fully determined by the concept set $\bm{c}$, or equivalently, $p(\bm{y}|\bm{c})\neq p(\bm{y}|\bm{c},\bm{x})$, which is typically the case in real-world scenarios [18].
-Concept Embedding Models (CEMs) [19] were proposed to mitigate the accuracy–interpretability trade-off in CBMs.
-However, this trade-off is fundamentally limited by the chosen concept set, and CEMs may even be less interpretable than standard CBMs, as their more entropic representations tend to amplify information leakage.
-Recently, this critique has been further formalized: Parisini et al.
-[20] proposed information-theoretic benchmarks for assessing interpretability and Makonnen et al.
-[21] introduced new metrics grounded in usable mutual information [22].
+Information leakage occurs when the learned representation $\bm{z}$ encodes information outside the concept set $\bm{c}$, reducing both interpretability and intervenability (Margeloiu et al., 2021; Mahinpei et al., 2021).
+This arises when the Markovian assumption fails—i.e., when the target $\bm{y}$ is not fully determined by the concept set $\bm{c}$, or equivalently, $p(\bm{y}|\bm{c})\neq p(\bm{y}|\bm{c},\bm{x})$, which is typically the case in real-world scenarios (Havasi et al., 2022).
+Concept Embedding Models (CEMs) (Espinosa Zarlenga et al., 2022) were proposed to mitigate the accuracy–interpretability trade-off in CBMs.
+However, this trade-off is fundamentally limited by the chosen concept set, and CEMs may even be less interpretable than standard CBMs, as their more entropic representations tend to amplify information leakage, a critique formalized in (Parisini et al., 2025).
 Havasi et al.
-[18] also introduced Hard Concept Bottleneck Models (HCBMs), which predict $\bm{y}$ from binarized concept predictions $\hat{c}_{j}$ rather than from $\bm{z}$ (see Appendix D), thereby imposing an ad-hoc Information Bottleneck.
-Extensions such as Autoregressive CBMs and Stochastic CBMs [18, 23] incorporate dependencies between concepts.
+(2022) also introduced Hard Concept Bottleneck Models (HCBMs), which predict $\bm{y}$ from binarized concept predictions $\hat{c}_{j}$ rather than from $\bm{z}$ (see Appendix D), thereby imposing an ad-hoc Information Bottleneck.
+Extensions such as Autoregressive CBMs (ARCBMs) and Stochastic CBMs (SCBMs)(Havasi et al., 2022; Vandenhirtz et al., 2024) incorporate dependencies between concepts.
+Energy-based CBMs (Xu et al., 2024a) replace the task head with an energy function that scores concept–label compatibility, enabling richer and more structured concept relationships.
+Beyond architectural changes, prior work has characterized leakage using information-theoretic quantities (Parisini et al., 2025; Makonnen et al., 2025).
+To the best of our knowledge, this is the first work that leverages Information Theory to (i) formally identify the underlying design flaw—CBMs require each $z_{j}$ to predict $c_{j}$ but do not restrict $z_{j}$ from encoding additional nuisance information—and (ii) introduce a principled, variational IB objective that directly constrains each latent variable to retain only concept-relevant information.
 Information Bottleneck in Representation Learning
-The Information Bottleneck (IB) [12] provides a principled way to balance preserving information about a factor with compressing the representation.
-In this framework, a representation is sufficient if it retains all the information about the variable of interest, and minimal if it contains only that information [24, 25, 26].
-Computation of these quantities is intractable, so variational methods have been developed to derive tractable evidence bounds [27, 28].
+The Information Bottleneck (IB) (Tishby et al., 2000) provides a principled way to balance preserving information about a factor with compressing the representation.
+In this framework, a representation is sufficient if it retains all the information about the variable of interest, and minimal if it contains only that information (Achille and Soatto, 2018a; b; Shwartz Ziv and LeCun, 2024).
+Computation of these quantities is intractable, so variational methods have been developed to derive tractable evidence bounds (Alemi et al., 2016; Fischer, 2020).
 4 Experiments
@@
 In contrast, we show that MCBMs successfully enforce a bottleneck, thereby mitigating the issues that arise in its absence.
-Further implementation details—including encoder architectures and training hyperparameters for each experiment—are provided in Appendix E and at https://github.com/antonioalmudevar/minimal_cbm.
+Further implementation details—including encoder architectures and training hyperparameters for each experiment—are provided in Appendix F.
 4.1 Do CBMs and MCBMs Leak Information?
@@
 We next examine whether $\bm{n}_{y}$ and $\bm{n}_{\bar{y}}$ are present in the representations across different CBM variants and datasets.
-For this purpose, we define the following task–concept configurations: (i) MPI3D[29], where $\bm{y}$ is the object shape, $\bm{n}_{y}$ the horizontal axis, $\bm{n}_{\bar{y}}$ the vertical axis, and $\bm{c}$ the remaining generative factors; (ii) Shapes3D[30], where $\bm{y}$ is the shape, $\bm{n}_{y}$ the floor color and wall color, $\bm{n}_{\bar{y}}$ the orientation, and $\bm{c}$ the remaining factors; (iii) CIFAR-10[31], where $\bm{y}$ is the standard classification task, $\bm{c}$ consists of 64 of the 143 attributes extracted by Oikarinen et al.
-[32] using GPT-3 [33], and $\bm{n}_{y}$ the remaining attributes, since all nuisances are correlated with $\bm{y}$; and (iv) CUB[34], where $\bm{y}$ is the bird species, $\bm{c}$ includes concepts from twelve randomly selected attribute groups, and $\bm{n}_{y}$ the attributes from the remaining 20 groups.
+For this purpose, we define the following task–concept configurations: (i) MPI3D(Gondal et al., 2019), where $\bm{y}$ is the object shape, $\bm{n}_{y}$ the horizontal axis, $\bm{n}_{\bar{y}}$ the vertical axis, and $\bm{c}$ the remaining generative factors; (ii) Shapes3D(Kim and Mnih, 2018), where $\bm{y}$ is the shape, $\bm{n}_{y}$ the floor color and wall color, $\bm{n}_{\bar{y}}$ the orientation, and $\bm{c}$ the remaining factors; (iii) CIFAR-10(Krizhevsky et al., 2009), where $\bm{y}$ is the standard classification task, $\bm{c}$ consists of 64 of the 143 attributes extracted by Oikarinen et al.
+(2023) using GPT-3 (Brown et al., 2020), and $\bm{n}_{y}$ the remaining attributes, since all nuisances are correlated with $\bm{y}$; (iv) CUB(Wah et al., 2011), where $\bm{y}$ is the bird species, $\bm{c}$ includes concepts from twelve randomly selected attribute groups, and $\bm{n}_{y}$ the attributes from the remaining 15 groups; and (v) AwA2(Xian et al., 2017), where $\bm{y}$ is the animal class, $\bm{c}$ consists of 20 of the 85 human-annotated attributes, and $\bm{n}_{y}$ the remaining attributes.
 While MCBMs natively support multiclass concepts, the other baselines in our study are limited to binary concepts.
 To ensure fairness, factors with $k$ classes are therefore represented as $k$ binary concepts.
-Table 3: Average value of URR for task-related nuisances $\bm{n}_{y}$.
-MPI3D Shapes3D CIFAR-10 CUB
-Vanilla $35.0\pm 1.9$ $45.5\pm 6.4$ $19.8\pm 0.7$ $3.8\pm 1.0$
-CBM [9] $28.1\pm 0.5$ $18.1\pm 2.9$ $18.5\pm 0.7$ $3.8\pm 0.8$
-CEM [19] $43.2\pm 5.2$ $15.8\pm 3.9$ $27.2\pm 0.8$ $3.9\pm 1.1$
-ARCBM [18] $28.2\pm 1.7$ $18.4\pm 2.1$ $18.2\pm 0.6$ $3.9\pm 0.9$
-SCBM [23] $24.3\pm 0.4$ $21.8\pm 1.4$ $18.3\pm 0.7$ $3.6\pm 0.8$
-MCBM (low $\gamma$) $10.7\pm 0.1$ $2.5\pm 0.1$ $18.0\pm 0.5$ $3.4\pm 0.9$
-MCBM (medium $\gamma$) $6.7\pm 0.2$ $0.2\pm 0.3$ $18.1\pm 0.5$ $2.8\pm 0.8$
-MCBM (high $\gamma$) $\bm{0.0\pm 0.0}$ $\bm{0.0\pm 0.0}$ $\bm{17.6\pm 0.5}$ $\bm{2.4\pm 1.0}$
 Task-related information leakage
@@
 In Table 3, we report the average value of $\frac{\hat{I}(N_{j};Z\mid C)}{H(N_{j})}$ across all $n_{j}\in\bm{n}_{y}$, which we call Uncertainty Reduction Ratio (URR).
-From these results, we conclude that: (i) CBMs tend to reduce nuisance information compared to VMs, though not consistently; (ii) CEMs generally preserve the largest amount of nuisance information, often exceeding even VMs; (iii) ARCBMs and HCBMs show no systematic advantage over CBMs in terms of nuisance removal; and (iv) MCBMs provide the strongest reduction of nuisance information, particularly as $\gamma$ increases, which enforces a stricter bottleneck.
+From these results, we conclude that: (i) CBMs tend to reduce nuisance information compared to VMs, though not consistently; (ii) CEMs and ECBMs generally preserve the largest amount of nuisance information, often exceeding even VMs; (iii) ARCBMs and HCBMs show no systematic advantage over CBMs in terms of nuisance removal; and (iv) MCBMs provide the strongest reduction of nuisance information, particularly as $\gamma$ increases, which enforces a stricter bottleneck.
+Table 3: Average value of URR for task-related nuisances $\bm{n}_{y}$.
+MPI3D Shapes3D CIFAR-10 CUB AwA2
+Vanilla $35.0\pm 1.9$ $45.5\pm 6.4$ $19.8\pm 0.7$ $3.8\pm 1.0$ $1.5\pm 0.3$
+CBM $28.1\pm 0.5$ $18.1\pm 2.9$ $18.5\pm 0.7$ $3.8\pm 0.8$ $1.4\pm 0.3$
+CEM $43.2\pm 5.2$ $15.8\pm 3.9$ $27.2\pm 0.8$ $3.9\pm 1.1$ $1.1\pm 0.5$
+ECBM $25.2\pm 3.0$ $47.1\pm 3.7$ $18.1\pm 0.5$ $4.5\pm 1.0$ $1.1\pm 0.3$
+ARCBM $28.2\pm 1.7$ $18.4\pm 2.1$ $18.2\pm 0.6$ $3.9\pm 0.9$ $1.6\pm 0.3$
+SCBM $24.3\pm 0.4$ $21.8\pm 1.4$ $18.3\pm 0.7$ $3.6\pm 0.8$ $1.2\pm 0.2$
+MCBM (low $\gamma$) $10.7\pm 0.1$ $2.5\pm 0.1$ $18.0\pm 0.5$ $3.4\pm 0.9$ $1.0\pm 0.4$
+MCBM (medium $\gamma$) $6.7\pm 0.2$ $0.2\pm 0.3$ $18.1\pm 0.5$ $2.8\pm 0.8$ $0.9\pm 0.4$
+MCBM (high $\gamma$) $\bm{0.0\pm 0.0}$ $\bm{0.0\pm 0.0}$ $\bm{17.6\pm 0.5}$ $\bm{2.4\pm 1.0}$ $0.7\pm 0.4$
 Task-unrelated information leakage
-Table 4: Average value of URR for task-unrelated nuisances.
+Table 4: Average value of URR for task- nuisances.
 MPI3D Shapes3D
@@
 CEM $15.5\pm 4.2$ $40.9\pm 1.8$
+ECBM $6.2\pm 1.4$ $46.4\pm 4.7$
 ARCBM $8.7\pm 1.4$ $26.6\pm 0.5$
@@
 MCBM (h $\gamma$) $\bm{0.0\pm 0.0}$ $\bm{0.0\pm 0.0}$
-Neural networks typically discard input information, as their layers are non-invertible [35, 36].
+Neural networks typically discard input information, as their layers are non-invertible (Tishby and Zaslavsky, 2015; Tschannen et al., 2019).
 Moreover, since $\bm{n}_{\bar{y}}$ is irrelevant for predicting $\bm{y}$, there is no incentive to retain it in the representation $\bm{z}$.
 One might therefore expect $\bm{z}$ to be free of $\bm{n}_{\bar{y}}$.
-However, prior work has shown that neural representations often preserve information not directly related to the task [24, 37].
+However, prior work has shown that neural representations often preserve information not directly related to the task (Achille and Soatto, 2018a; Arjovsky et al., 2019).
 To examine this, we analyze whether $\bm{n}_{\bar{y}}$ is present in $\bm{z}$ for MPI3D and Shapes3D—the only settings where $\bm{n}_{\bar{y}}$ is non-empty.
 Table 4 reports URR values for the nuisance variables in $\bm{n}_{\bar{y}}$ in.
-We find that: (i) as in Table 3, CEMs retain more nuisance information than even VMs; and (ii) MCBMs are the only models that consistently eliminate nuisances across all values of $\gamma$, as expected: with no incentive to preserve $\bm{n}_{\bar{y}}$ and an explicit penalty for doing so, such information is naturally discarded.
+We find that: (i) as in Table 3, CEMs and ECBMs can retain more nuisance information than even VMs; and (ii) MCBMs are the only models that consistently eliminate nuisances across all values of $\gamma$, as expected: with no incentive to preserve $\bm{n}_{\bar{y}}$ and an explicit penalty for doing so, such information is naturally discarded.
 Are concepts removed to a greater extent in MCBMs?
 Table 5: Average concepts accuracy
-CIFAR-10 CUB
-CBM $84.8\pm 0.2$ $96.3\pm 0.1$
-CEM $84.8\pm 0.2$ $96.3\pm 0.1$
-ARCBM $84.3\pm 0.2$ $96.2\pm 0.1$
-SCBM $84.3\pm 0.2$ $\bm{96.5\pm 0.1}$
-MCBM (l $\gamma$) $\bm{84.9\pm 0.1}$ $96.3\pm 0.2$
-MCBM (m $\gamma$) $\bm{84.9\pm 0.2}$ $96.1\pm 0.1$
-MCBM (h $\gamma$) $84.8\pm 0.1$ $95.8\pm 0.3$
+CIFAR-10 CUB AwA2
+CBM $84.8\pm 0.2$ $96.3\pm 0.1$ $98.1\pm 0.1$
+CEM $84.8\pm 0.2$ $96.3\pm 0.1$ $98.0\pm 0.1$
+ECBM $84.6\pm 0.2$ $96.1\pm 0.4$ $98.0\pm 0.1$
+ARCBM $84.3\pm 0.2$ $96.2\pm 0.1$ $97.9\pm 0.1$
+SCBM $84.3\pm 0.2$ $\bm{96.5\pm 0.1}$ $\bm{98.4\pm 0.1}$
+MCBM (l $\gamma$) $\bm{84.9\pm 0.1}$ $96.3\pm 0.2$ $98.0\pm 0.2$
+MCBM (m $\gamma$) $\bm{84.9\pm 0.2}$ $96.1\pm 0.1$ $97.9\pm 0.1$
+MCBM (h $\gamma$) $84.8\pm 0.1$ $95.8\pm 0.3$ $97.6\pm 0.2$
 One might worry that removing nuisance information could also inadvertently eliminate information about the concepts.
@@
 How does this affect task performance?
+As previously discussed, when $\bm{n}_{y}$ is non-empty, restricting $\bm{z}$ to contain only information about $\bm{c}$ should reduce performance on $\bm{y}$: if $\bm{c}$ alone is insufficient to solve $\bm{y}$, then $\bm{z}$ cannot achieve perfect accuracy.
+In Table 6, we observe the following: (i) CBMs, CEMs and ECBMs reach task accuracy comparable to (or even higher than) VMs, indicating that they also rely on $\bm{n}_{y}$ to predict $\bm{y}$; (ii) ARCBMs and SCBMs achieve lower task accuracy, as they impose a bottleneck after the representations (see Appendix D); and (iii) MCBMs show decreasing task accuracy as $\gamma$ increases, reflecting the stricter bottleneck applied to the representations.
+Importantly, this reduction in task accuracy should not be viewed negatively: since all models achieve similar concept accuracy (see Table 5), it indicates that predictions rely less on nuisance information.
 Table 6: Task accuracy
-MPI3D Shapes3D CIFAR-10 CUB
-Vanilla $\pm 0.0 100.0±0.010$$\pm 0.0$ $2.1\pm 0.6 77$$7.4\pm 0.3$
-CBM $99.9\pm 0.0$ $100.0\pm 0.0$ $3.8\pm 0.1 77$$7.6\pm 0.2$
-CEM $\pm 0.0 100.0±0.010$$\pm 0.0$ $3.1\pm 0.3 77$$7.5\pm 0.3$
-ARCBM $24.2\pm 0.6$ $29.7\pm 0.4$ $68.9\pm 0.3$ $75.5\pm 0.8$
-SCBM $24.2\pm 0.5$ $29.7\pm 0.3$ $62.3\pm 0.0$ $73.5\pm 0.4$
-MCBM (l $\gamma$) $92.7\pm 1.1$ $100.0\pm 0.0$ $72.4\pm 0.5$ $78.3\pm 0.6$
-MCBM (m $\gamma$) $46.0\pm 0.6$ $32.2\pm 1.5$ $0.8\pm 0.2 77$$7.4\pm 0.4$
-MCBM (h $\gamma$) $24.9\pm 0.3$ $30.1\pm 0.2$ $70.5\pm 0.6$ $73.5\pm 1.5$
-As previously discussed, when $\bm{n}_{y}$ is non-empty, restricting $\bm{z}$ to contain only information about $\bm{c}$ should reduce performance on $\bm{y}$: if $\bm{c}$ alone is insufficient to solve $\bm{y}$, then $\bm{z}$ cannot achieve perfect accuracy.
-In Table 6, we observe the following: (i) CBMs and CEMs reach task accuracy comparable to (or even higher than) VMs, indicating that they also rely on $\bm{n}_{y}$ to predict $\bm{y}$; (ii) ARCBMs and SCBMs achieve lower task accuracy, as they impose a bottleneck after the representations (see Appendix D); and (iii) MCBMs show decreasing task accuracy as $\gamma$ increases, reflecting the stricter bottleneck applied to the representations.
-Importantly, this reduction in task accuracy should not be viewed negatively: since all models achieve similar concept accuracy (see Table 5), it indicates that predictions rely less on nuisance information.
+MPI3D Shapes3D CIFAR-10 CUB AwA2
+Vanilla $\pm 0.0 100.0±0.010$$\pm 0.0$ $2.1\pm 0.6 77$$7.4\pm 0.3$ $91.6\pm 0.4$
+CBM $99.9\pm 0.0$ $100.0\pm 0.0$ $3.8\pm 0.1 77$$7.6\pm 0.2$ $91.1\pm 0.4$
+CEM $\pm 0.0 100.0±0.010$$\pm 0.0$ $3.1\pm 0.3 77$$7.5\pm 0.3$ $90.7\pm 0.5$
+ECBM $99.9\pm 0.0$ $100.0\pm 0.0$ $4.4\pm 0.4 77$$7.1\pm 0.5$ $92.1\pm 0.3$
+ARCBM $24.2\pm 0.6$ $29.7\pm 0.4$ $68.9\pm 0.3$ $75.5\pm 0.8$ $89.9\pm 0.3$
+SCBM $24.2\pm 0.5$ $29.7\pm 0.3$ $62.3\pm 0.0$ $73.5\pm 0.4$ $91.8\pm 0.1$
+MCBM (l $\gamma$) $92.7\pm 1.1$ $100.0\pm 0.0$ $72.4\pm 0.5$ $78.3\pm 0.6$ $90.6\pm 0.4$
+MCBM (m $\gamma$) $46.0\pm 0.6$ $32.2\pm 1.5$ $0.8\pm 0.2 77$$7.4\pm 0.4$ $90.1\pm 0.2$
+MCBM (h $\gamma$) $24.9\pm 0.3$ $30.1\pm 0.2$ $70.5\pm 0.6$ $73.5\pm 1.5$ $88.7\pm 0.3$
 4.2 Do MCBMs Yield More Interpretable Representations?
-Figure 4: CKA (left) and Disentanglement (right).
-One of the key properties often attributed to CBMs is that their internal representations align with human-interpretable concepts.
+One of the key properties often attributed to CBMs is that their internal representations align with human-interpretable concepts (Debole et al., 2025).
 While this generally holds, we show that MCBMs yield representations that are even more interpretable.
-To support this claim, we employ two metrics: (i) Centered Kernel Alignment(CKA) [38, 39, 40], which measures the similarity between learned representations and concept labels (encoded as one-hot vectors); and (ii) Disentanglement[41], which assesses whether each dimension of the representation $z_{j}$ captures at most one concept $c_{j}$.
-These metrics capture interpretability from complementary perspectives: (i) alignment indicates whether the information is organized in a concept-aware fashion; and (ii) disentanglement reflects the extent to which individual concepts are independently encoded in the representations.
-As shown in Figure 4: (i) CBMs, CEMs, ARCBMs, and HCBMs achieve stronger alignment with concepts than Vanilla Models, yet they do not consistently improve disentanglement; and (ii) MCBMs consistently improve both alignment and disentanglement, particularly as $\gamma$ increases, reflecting the removal of additional nuisances.
-These results suggest that explicitly eliminating nuisances leads to more interpretable representations.
+To support this claim, we employ two metrics: (i) Centered Kernel Alignment(CKA) (Cristianini et al., 2001; Cortes et al., 2012; Kornblith et al., 2019), which measures the similarity between learned representations and concept labels (encoded as one-hot vectors); (ii) Disentanglement(Eastwood and Williams, 2018), which assesses whether each dimension of the representation $z_{j}$ captures at most one concept $c_{j}$; and (iii) Oracle Information Score (OIS) (Zarlenga et al., 2023), which dextends beyond standard disentanglement by accounting for concept dependencies, requiring that correlations between $z_{j}$ and $z_{k}$ do not exceed those between $c_{j}$ and $c_{k}$.
+These metrics capture interpretability from complementary perspectives: (i) CKAindicates whether the information is organized in a concept-aware fashion; and (ii) Disentanglement and OIS reflect the extent to which individual concepts are independently encoded in the representations.
+As shown in Figure 4: As shown in Figure 4: (i) CBMs, CEMs, AR-CBMs, and HCBMs generally achieve higher CKA with concepts than Vanilla Models, but they do not consistently improve Disentanglement or OIS; (ii) ECBMs do not, in general, produce more interpretable representations than VMs; and (iii) MCBMs consistently improve CKA, Disentanglement, and OIS, with larger gains as $\gamma$ increases, reflecting the removal of additional nuisances.
+These results indicate that explicitly removing nuisances leads to more interpretable representations.
+To analyze this systematically, Table 7 reports the rank correlations between the metrics in Figure 4 and the URR values in Table 3 across datasets.
+We observe that CKA and OIS are strong and consistent predictors of nuisance information, whereas disentanglement is noticeably weaker.
+This is particularly helpful for tuning $\gamma$: since $n_{y}$ is unobserved in real scenarios, URR cannot be computed as in Table 3.
+As a workaround, these metrics—computed solely from $\bm{c}$ and $\bm{z}$—can serve as practical proxies for selecting the trade-off between task accuracy and leakage.
+Figure 4: CKA (left, ↑), Disentanglement (middle, ↑) and OIS (right, ↓).
+Table 7: Rank Correlation (p-value) between each metric and URR for $n_{y}$ across datasets
+MPI3D Shapes3D CIFAR-10 CUB AwA2
+CKA .95(<.001)-.95\,(<\!.001) .91(<.001)-.91\,(<\!.001) .71(.04)-.71\,(.04) .77(.02)-.77\,(.02) .75(.04)-.75\,(.04)
+Disentanglement .81(.01)-.81\,(.01) .57(.13)-.57\,(.13) .43(.29)-.43\,(.29) .69(.05)-.69\,(.05) .99(<.001)-.99\,(<\!.001)
+OIS .90(.002).90\,(.002) .81(.01).81\,(.01) .86(.006).86\,(.006) .89(.002).89\,(.002) .66(.07).66\,(.07)
 4.3 Do MCBMs Enable More Reliable Interventions?
-(a) CIFAR-10
-(b) CUB
-Figure 5: Error (y-axis) versus percentage of concepts intervened (x-axis) across different models.
 Another key property often attributed to CBMs is their capacity to support interventions.
@@
 To evaluate intervention reliability across models, we adopt the standardized protocol of Koh et al.
-[9], which tracks how prediction error changes as the number of intervened concepts increases.
+(2020), which tracks how prediction error changes as the number of intervened concepts increases.
 Following Shin et al.
-[42], we intervene on the concepts predicted with the lowest confidence.
-Results for CIFAR-10 and CUB are shown in Figure 5 (note that interventions have minimal effect on MPI3D and Shapes3D, since—as shown in Table 6—the concepts are not strong predictors of the task).
-From these results, four main observations emerge: (i) CBMs may even increase error when multiple concepts are intervened—an effect of nuisance information leaking into the representation due to the absence of a proper bottleneck; (ii) ARCBMs and SCBMs mitigate this by applying an Information Bottleneck after the representations, which improves intervention effectiveness while leaving interpretability unchanged (Figure 4); (iii) in MCBMs, intervention gains remain largely invariant to $\gamma$, as indicated by the nearly parallel trends across its different values; and (iv) at low and medium $\gamma$, MCBMs deliver the strongest intervention performance, especially when fewer than 100% of the concepts are intervened, clearly outperforming ARCBMs and SCBMs.
+(2023), we examine the effect of interventions using two complementary policies on CIFAR-10, CUB, and AwA2: (i) intervening on the concepts with the lowest predicted confidence (results in Figure 5); and (ii) intervening on randomly selected concepts (results in Appendix E).
+For MPI3D and Shapes3D, interventions yield only minor changes in performance, which aligns with Table 6 showing that their concepts are relatively weak predictors of the target.
+Across datasets where concepts are informative, both intervention policies lead to consistent and converging conclusions: (i) CBMs may even increase error when multiple concepts are intervened—an effect of nuisance information leaking into the representation due to the absence of a proper bottleneck; (ii) ARCBMs and SCBMs mitigate this by applying an Information Bottleneck after the representations, which improves intervention effectiveness while leaving interpretability unchanged (Figure 4); (iii) in MCBMs, intervention gains remain mostly insensitive to $\gamma$ when only a small or moderate number of concepts are intervened—evidenced by the nearly parallel curves across $\gamma$ values—but increase with higher $\gamma$ when more concepts are intervened, as reflected in the steeper negative slopes; (iv) at low and medium $\gamma$, MCBMs deliver the strongest intervention performance, especially when fewer than 100% of the concepts are intervened, clearly outperforming ARCBMs and SCBMs; and (v) for AwA2—where all models exhibit low levels of nuisance retention (Table 3)—intervention performance is nearly identical across methods.
+(a) CIFAR-10
+(b) CUB
+(c) AwA2
+Figure 5: Error (y-axis) versus percentage of concepts intervened (x-axis) across different models.
 5 Other Fundamental Theoretical Flaws of CBMs
@@
 One-vs-Rest Limitations
-Figure 6: CBM (top) versus MCBM (bottom): class boundaries (left), highest predicted likelihood (middle), and second-highest predicted likelihood (right).
-One-vs-Rest strategies exhibit several limitations: (i) individual binary classifiers tend to be biased toward the negative class, and (ii) the predicted probabilities across classifiers are typically uncalibrated [43].
+One-vs-Rest strategies exhibit several limitations: (i) individual binary classifiers tend to be biased toward the negative class, and (ii) the predicted probabilities across classifiers are typically uncalibrated (Bishop, 2006).
 To illustrate these issues, we design an experiment where the concepts are defined based on a four-class spiral dataset with imbalanced class distributions.
 We train: (i) a CBM with One-vs-Rest binarization, and (ii) an MCBM modeling concepts directly as multiclass variables.
+Figure 6: CBM (top) versus MCBM (bottom): class boundaries (left), highest predicted likelihood (middle), and second-highest predicted likelihood (right).
 The results reveal three major shortcomings of CBMs trained with the One-vs-Rest strategy.
@@
 […]
-6 Conclusions
+6 Conclusions, Limitations and Future Directions
 In this paper, we argue that—contrary to common belief—Concept Bottleneck Models (CBMs) do not enforce a true bottleneck: although representations are encouraged to retain concept-related information, they are not constrained to discard nuisance information.
@@
 Finally, we highlight fundamental flaws in the intervention process of CBMs, which MCBMs overcome due to their principled design.
-Acknowledgments and Disclosure of Funding
-This work has received funding from the European Union’s Horizon 2020 research and innovation programme under the Marie Skłodowska-Curie grant agreement No 101007666, MCIN/AEI/10.13039/501100011033 under Grant PID2021-126061OB-C44, and the Government of Aragón (Grant Group T36 23R).
+Regarding the limitations of this work, we highlight the following: (i) MCBMs introduce a new hyperparameter, $\gamma$, which must be tuned to balance predictive accuracy and interpretability; (ii) the representation head $g_{\phi}^{z}$ adds a small number of parameters, though this overhead is negligible relative to the backbone; and (iii) while MCBMs consistently reduce nuisance information, complete removal remains challenging for high-variance datasets, suggesting that performance depends on the expressive capacity of the backbone architecture.
+As for future directions, promising avenues include: (i) extending the model with an auxiliary latent variable $z_{m+1}$ appended to $\bm{z}$ to explicitly capture task-relevant nuisance factors $n_{y}$, allowing $z_{1},\dots,z_{m}$ to remain strictly interpretable while still retaining task-useful information in the full representation $\bm{z}$; and (ii) studying how the choice of prior distributions $q_{\phi}(z_{j}\mid c_{j})$ affects representation quality and evaluation metrics such as disentanglement, alignment, and concept leakage.
+Acknowledgments
+This work has received funding from the European Union’s Horizon 2020 research and innovation programme under the Marie Skłodowska-Curie grant agreement No 101007666, MCIN/AEI/10.13039/501100011033 under Grant PID2024-155948OB-C53, and the Government of Aragón (Grant Group T36 23R).
+Code Availability
+An implementation accompanying this work is available at https://github.com/antonioalmudevar/minimal_cbm
 Appendix A Complete Graphical Models
@@
 Appendix D Hard Concept Bottleneck Models
-As discussed in Section 3, one of the most successful approaches to mitigating information leakage is the family of Hard Concept Bottleneck Models (HCBMs) [18].
+As discussed in Section 3, one of the most successful approaches to mitigating information leakage is the family of Hard Concept Bottleneck Models (HCBMs) (Havasi et al., 2022).
 Unlike MCBMs, which enforce an Information Bottleneck (IB) directly at the representation level $\bm{z}$, HCBMs operate by pruning the information contained in $\bm{z}$ before producing the task prediction $\hat{\bm{y}}$.
@@
 Figure 9: Graphical models of HCBMs with two concepts and two-dimensional representations.
-Hard Concept Bottleneck Models obtain a binarized version $\hat{c}^{b}_{j}$ of each predicted concept $\hat{c}_{j}$ through the binarizing head $\color[rgb]{1,0,1}p(\hat{c}^{b}_{j}\mid\hat{c}j)$ (fuchsia arrows).
-Unlike the models in Figure 2, HCBMs predict the task output $\hat{\bm{y}}$ from the binarized concepts $\hat{c}^{b}_{j}$ using the new task head $\color[rgb]{0,1,0}q_{\phi}(\hat{\bm{y}}\mid\bm{z})$ (green arrows).
-Appendix E Experiments Details
-E.1 Hyperparameters for Section 4
+Hard Concept Bottleneck Models obtain a binarized version $\hat{c}^{b}_{j}$ of each predicted concept $\hat{c}_{j}$ through the binarizing head $\color[rgb]{1,0,1}p(\hat{c}^{b}_{j}\mid\hat{c}_{j})$ (fuchsia arrows).
+Unlike the models in Figure 2, HCBMs predict the task output $\hat{\bm{y}}$ from the binarized concepts $\hat{c}^{b}_{j}$ using the new task head $\color[rgb]{0,1,0}q_{\phi}(\hat{\bm{y}}\mid\hat{\bm{c}}^{b})$ (green arrows).
+Appendix E Other Intervention Procedures
+As detailed in Section 4.3, we further examine how performance evolves as we intervene on an increasing number of randomly selected concepts, with the resulting curves shown in Figure 10.
+The overall behavior is consistent with the trends observed in Figure 5, highlighting similar model characteristics.
+As expected, the results exhibit somewhat higher variance across seeds, since each run intervenes on a different subset of concepts.
+(a) CIFAR-10
+(b) CUB
+(c) AwA2
+Figure 10: Error (y-axis) versus percentage of concepts intervened (x-axis) across different models for randomly selected concepts.
+Appendix F Experiments Details
+F.1 Hyperparameters for Section 4
 Table 8: Hyperparameters for Section 4.
@@
 We emphasize that CBMs are, by design, restricted to use invertible $g^{c}_{\phi}$ to enable intervention procedures.
-MPI3D Shapes3D CIFAR-10 CUB
-$f_{\theta}$ architecture ResNet20 ResNet20 2 conv. layers InceptionV3
-$f_{\theta}$ pretraining None None None ImageNet
-$g^{y}_{\phi}$ hidden layers 64 64 64 256
-low $\gamma$ 1 1 0.1 0.05
-medium $\gamma$ 3 3 0.3 0.1
-high $\gamma$ 5 5 0.5 0.3
-number of epochs 50 50 200 250
-batch size 128 128 128 128
-optimizer SGD SGD Adam SGD
-learning rate $6\times 10^{-3}$ $6\times 10^{-3}$ $1\times 10^{-4}$ $2\times 10^{-2}$
-momentum 0.9 0.9 0. 0.9
-weight decay $4\times 10^{-5}$ $4\times 10^{-5}$ $4\times 10^{-5}$ $4\times 10^{-5}$
-scheduler Step Step Step Step
-step size (epochs) 20 20 80 100
-scheduler $\gamma$ 0.1 0.1 0.1 0.1
-E.2 Datasets
+MPI3D Shapes3D CIFAR-10 CUB AwA2
+$f_{\theta}$ architecture ResNet20 ResNet20 2 conv. layers InceptionV3 ResNet-50
+$f_{\theta}$ pretraining None None None ImageNet ImageNet
+$g^{y}_{\phi}$ hidden layers 64 64 64 256 256
+$g^{c}_{\phi}$ hidden layers None None None None None
+$g^{z}_{\phi}$ hidden layers [^1] Beyond the choice of $\gamma$, MCBMs introduce an additional design decision: the architecture of $g_{\phi}^{z}(c_{j})$.
+In all our experiments, we implement this module as a small MLP with a single hidden layer of size 3, adding only 8 parameters per concept.
+Since concept sets typically contain at most 200 concepts, this corresponds to roughly 1600 additional parameters—negligible compared to the size of standard neural encoders $f_{\theta}$.
+3 3 3 3 3
+low $\gamma$ 1 1 0.1 0.05 0.05
+medium $\gamma$ 3 3 0.3 0.1 0.1
+high $\gamma$ 5 5 0.5 0.3 0.3
+number of epochs 50 50 200 250 120
+batch size 128 128 128 128 128
+optimizer SGD SGD Adam SGD SGD
+learning rate $6\times 10^{-3}$ $6\times 10^{-3}$ $1\times 10^{-4}$ $2\times 10^{-2}$ $2\times 10^{-2}$
+momentum 0.9 0.9 0. 0.9 0.9
+weight decay $4\times 10^{-5}$ $4\times 10^{-5}$ $4\times 10^{-5}$ $4\times 10^{-5}$ $4\times 10^{-5}$
+scheduler Step Step Step Step Step
+step size (epochs) 20 20 80 100 50
+scheduler $\gamma$ 0.1 0.1 0.1 0.1 0.1
+F.2 Datasets
 MPI3D
@@
 The dataset is split into 50,000 training and 10,000 test images, with balanced class distributions.
-To reduce the need for manual concept annotations, the concepts are synthetically derived following the methodology of [23].
-A total of 143 attributes are extracted using GPT-3 [33]; 64 form the concept set $\bm{c}$, while the rest define the nuisance set $\bm{n}_{y}$.
-Binary values are obtained with the CLIP model [44] by comparing the similarity of each image to the embedding of an attribute and to its negative counterpart.
+To reduce the need for manual concept annotations, the concepts are synthetically derived following the methodology of (Vandenhirtz et al., 2024).
+A total of 143 attributes are extracted using GPT-3 (Brown et al., 2020); 64 form the concept set $\bm{c}$, while the rest define the nuisance set $\bm{n}_{y}$.
+Binary values are obtained with the CLIP model (Radford et al., 2021) by comparing the similarity of each image to the embedding of an attribute and to its negative counterpart.
 CUB
@@
 Following the approach of Koh et al.
-[9], we retain only the attributes that are present in at least 10 species (based on majority voting), resulting in a filtered set of 112 attributes.
+(2020), we retain only the attributes that are present in at least 10 species (based on majority voting), resulting in a filtered set of 112 attributes.
 These attributes are grouped into 27 semantic clusters, where each group is defined by a common prefix in the attribute names.
 In our setup, the task variable $\bm{y}$ is to classify the bird species.
-The concept set $\bm{c}$ consists of the attributes belonging to 12 randomly selected groups (per run), while the nuisance set $\bm{n}y$ includes the attributes from the remaining 20 groups.
+The concept set $\bm{c}$ consists of the attributes belonging to 12 randomly selected groups (per run), while the nuisance set $\bm{n}_{y}$ includes the attributes from the remaining 15 groups.
 Since most attributes exhibit some correlation with the classification task, we set $\bm{n}_{\bar{y}}$ to the empty set.
+AwA2
+The Animals with Attributes 2 (AwA2) dataset (Xian et al., 2017) contains 37,322 images of 50 animal classes annotated with 85 human-defined attributes describing appearance, behavior, and habitat.
+In our setup, the task variable $\bm{y}$ is the animal class.
+To construct the concept and nuisance partitions, we retain a subset of 20 attributes—covering fundamental appearance and morphology features—as the concept set $\bm{c}$, while the remaining 65 attributes form the nuisance set $\bm{n}_{y}$.
+Since nearly all attributes exhibit some degree of correlation with the class label, we set $\bm{n}_{\bar{y}}$ to the empty set.
+Following the preprocessing in Xian et al.
+(2017), we binarize continuous attributes using a threshold of 0.5.